### v1

# Transformers as Unsupervised Learning Algorithms: A study on Gaussian Mixtures

Abstract
The transformer architecture has demonstrated remarkable capabilities in modern artificial intelligence, among which the capability of implicitly learning an internal model during inference time is widely believed to play a key role in the under standing of pre-trained large language models. However, most recent works have been focusing on studying supervised learning topics such as in-context learning, leaving the field of unsupervised learning largely unexplored. This paper investigates the capabilities of transformers in solving Gaussian Mixture Models (GMMs), a fundamental unsupervised learning problem through the lens of statistical estimation. We propose a transformer-based learning framework called TGMM that simultaneously learns to solve multiple GMM tasks using a shared transformer backbone. The learned models are empirically demonstrated to effectively mitigate the limitations of classical methods such as Expectation-Maximization (EM) or spectral algorithms, at the same time exhibit reasonable robustness to distribution shifts. Theoretically, we prove that transformers can approximate both the EM algorithm and a core component of spectral methods (cubic tensor power iterations). These results bridge the gap between practical success and theoretical understanding, positioning transformers as versatile tools for unsupervised learning.  [^1] Code available at https://github.com/Rorschach1989/transformer-for-gmm
[…]
1 Introduction
Large Language Models (LLMs) have achieved remarkable success across various tasks in recent years. Transformers attention-is-all-you-need , the dominant architecture in modern LLMs BrownGPT3 , outperform many other neural network models in efficiency and scalability. Beyond language tasks, transformers have also demonstrated strong performance in other domains, such as computer vision survey-vision-TF ; TF-in-vision and reinforcement learning li2023TF-RL . Given their practical success, understanding the mechanisms behind transformers has attracted growing research interest. Existing studies often treat transformers as algorithmic toolboxes, investigating their ability to implement diverse algorithms (von-oswald23a, ; bai2023tfstats, ; lin2024transformers, ; giannou2025how, ; teh2025empiricalbayestransformers, )–a perspective linked to meta-learning hospedales2021meta .
However, most research has focused on supervised learning settings, such as regression bai2023tfstats and classification giannou2025how , leaving the unsupervised learning paradigm relatively unexplored. Since transformer models are typically trained in a supervised manner, unsupervised learning poses inherent challenges for transformers due to the absence of labeled data. Moreover, given the abundance of unlabeled data in real-world scenarios, investigating the mechanisms of transformers in unsupervised learning holds significant implications for practical applications. The Gaussian mixture model (GMM) represents one of the most fundamental unsupervised learning tasks in statistics, with a rich historical background(GMM1969, ; Aitkin01081980, ) and ongoing research interest(gmm2021ACM-TOMM, ; dcgmm-nips-2021, ; loffler2021optimality, ; ndaoud2022sharp, ; gribonval2021statistical, ; yu2021scgmai, ). Two primary algorithmic approaches are existing for solving GMM problems: (1) likelihood-based methods employing the Expectation-Maximization (EM) algorithm(EM1977, ; BinYu-EM-2017, ), and (2) moment-based methods utilizing spectral algorithms(10.1145LearningGMM, ; JMLR:v15:anandkumar14b, ). However, both algorithms have inherent limitations. The EM algorithm is prone to convergence at local optima and is highly sensitive to initialization moitra2018algorithmic ; chijinEM . In contrast, while the spectral method is independent of initialization, it requires the number of components to be smaller than the data’s dimensionality—an assumption that restricts its applicability to problems involving many components in low-dimensional GMMs10.1145LearningGMM . In this work, we explore transformers for GMM parameter estimation. We empirically find that meta-trained transformers exhibit strong performance on GMM tasks without the aforementioned limitations. Notably, we construct transformer-based solvers that efficiently solve GMMs with varying component counts simultaneously. The experimental phenomena are further backed up by novel theoretical establishments: We prove that transformers can effectively learn GMMs with different components by approximating both the EM algorithm and a key component of spectral methods on GMM tasks.
Our contributions are mainly summarized as follows.
  • We propose the TGMM framework that utilizes transformers to solve multiple GMM tasks with varying numbers of components simultaneously during inference time. Through extensive experimentation, the learned TGMM model is demonstrated to achieve competitive and robust performance over synthetic GMM tasks. Notably, TGMM outperforms the popular EM algorithm in terms of estimation quality, and approximately matches the strong performance of spectral methods while enjoying better flexibility.
  • We establish theoretical foundations by proving that transformers can approximate both the EM algorithm and a key component of spectral methods. Our approximation of the EM algorithm fundamentally leverages the weighted averaging property inherent in softmax attention, enabling simultaneous approximation of both the E and M steps. Notably, our approximation results also hold across varying dimensions and mixture components in GMM.
  • We prove that ReLU-activated transformers can exactly approximate cubic tensor power iterations – a crucial component of spectral algorithms for GMM. This result stems from a novel insight about attention heads: they can perform computations along dimensions beyond the query-key-value framework. To the best of our knowledge, this is the first theoretical demonstration of transformers’ capacity for high-order tensor calculations.
  • Our work is among the earliest studies to investigate the mechanism of transformers in unsupervised learning settings. Our findings enhance the theoretical understanding of transformers and offer valuable insights for future research.
The paper is organized as follows. In Section 2, some background knowledge is introduced. In Section 3, we present the experimental details and findings. The theoretical results are proposed in Section 4, and some discussions are given in Section 5. The proofs and additional experimental results are given in the appendix.
Related works. Recent research has explored the mechanisms by which transformers can implement various supervised learning algorithms. For instance, akyrek2023what , von-oswald23a , and bai2023tfstats demonstrate that transformers can perform gradient descent for linear regression problems in-context. lin2024transformers shows that transformers are capable of implementing Upper Confidence Bound (UCB) algorithms, as well as other classical algorithms in reinforcement learning tasks. giannou2025how reveals that transformers can execute in-context Newton’s method for logistic regression problems. teh2025empiricalbayestransformers illustrates that transformers can approximate Robbins’ estimator and solve Naive Bayes problems. kim2024transformers studies the minimax optimality of transformers on nonparametric regression.
Comparison with prior theoretical works in unsupervised learning setting. Several recent studies have investigated the mechanisms of transformer-based models in mixture model settings he2025learningspectralmethodstransformers ; jin2025incontextlearningmixturelinear ; he2025transformersversusemalgorithm . Among these, he2025learningspectralmethodstransformers establishes that transformers can implement Principal Component Analysis (PCA) and leverages this to GMM clustering. However, their analysis is limited to the two-component case, restricting its broader applicability. The paper jin2025incontextlearningmixturelinear investigates the in-context learning capabilities of transformers for mixture linear models, a setting that differs from ours. Furthermore, their approximation construction of the transformer is limited to two-component GMMs, leaving the general case unaddressed. While they assume ReLU as the activation function–contrary to the conventional choice of softmax–their theoretical proofs rely on a key lemma from prior workpathak2024transformers that assumes softmax activation, thereby introducing an inconsistency in their assumptions. The paper he2025transformersversusemalgorithm studies the performance of transformers on multi-class GMM clustering, a setting closely related to ours. However, our work focuses on parameter estimation rather than clustering. From a technical standpoint, their approximation bound scales exponentially with the dimension $d$, and their analysis requires the number of attention heads $M$ to grow to infinity to achieve a meaningful bound. In contrast, our results are significantly stronger: our bounds depend polynomially on $d$, and we only require $M=O(1)$. From an empirical perspective, their experiments are conducted on a small-scale transformer, which fails to validate their theoretical claims.
Notations. We introduce the following notations. Let $[n]\mathrel{\mathop{:}}={\{1,2,\cdots,n\}}$. All vectors are represented as column vectors unless otherwise specified. For a vector $v\in\mathbb{R}^{d}$, we denote $\left\|{v}\right\|$ as its Euclidean norm. For two sequences $a_{n}$ and $b_{n}$ indexed by $n$, we denote $a_{n}=O(b_{n})$ if there exists a universal constant $C$ such that $a_{n}\leq Cb_{n}$ for sufficiently large $n$.
2 Methodology
2.1 Preliminaries
The Gaussian mixture model (GMM) is a cornerstone of unsupervised learning in statistics, with deep historical roots and enduring relevance in modern research. Since its early formalizations (GMM1969, ; Aitkin01081980, ), GMM has remained a fundamental tool for clustering and density estimation, widely applied across diverse domains. Recent advances have further explored the theoretical foundations of Gaussian Mixture Models (GMMs) (loffler2021optimality, ; ndaoud2022sharp, ; gribonval2021statistical, ), extended their applications in incomplete data settings (gmm2021ACM-TOMM, ), and integrated them with deep learning frameworks (dcgmm-nips-2021, ; yu2021scgmai, ). Due to their versatility and interpretability, GMMs remain indispensable in unsupervised learning, effectively bridging classical statistical principles with modern machine learning paradigms. We consider the (unit-variance) isotropic Gaussian Mixture Model with $K$ components, with its probability density function as
$p(x|{\boldsymbol{\theta}})=\sum_{k=1}^{K}\pi_{k}\phi(x;\mu_{k})~{},$ (1)
where $\phi(x;\mu)$ is the standard Gaussian kernel, i.e. $\phi(x;\mu)=\frac{1}{(2\pi)^{d/2}}\exp\left(-\frac{1}{2}(x-\mu)^{\top}(x-\mu)\right)$. The parameter ${\boldsymbol{\theta}}$ is defined as ${\boldsymbol{\theta}}={\boldsymbol{\pi}}\cup{\boldsymbol{\mu}}$, where ${\boldsymbol{\pi}}\mathrel{\mathop{:}}={\{\pi_{1},\pi_{2},\cdots,\pi_{K}\}}$, $\pi_{k}\in\mathbb{R}$ and ${\boldsymbol{\mu}}={\{\mu_{1},\mu_{2},\cdots,\mu_{K}\}},\mu_{k}\in\mathbb{R}^{d}$, $k\in[K]$. We take $N$ samples ${\mathbf{X}}={\{X_{i}\}}_{i\in[N]}$ from model Eq. 1. ${\{X_{i}\}}_{i\in[N]}$ can be also rewritten as
$X_{i}=\mu_{y_{i}}+Z_{i},$
where ${\{y_{i}\}}_{i\in[N]}$ are iid discrete random variables with ${\mathbb{P}}{\left(y=k\right)}=\pi_{k}$ for $k\in[K]$ and ${\{Z_{i}\}}_{i\in[N]}$ are iid standard Gaussian random vector in $\mathbb{R}^{d}$.
The EM algorithm EM1977 remains the most widely used approach for GMM parameter estimation. Due to space constraints, we propose the algorithm in Appendix A. Alternatively, the spectral algorithm (10.1145LearningGMM, ) offers an efficient moment-based approach that estimates parameters through low-order observable moments. A key component of this method is cubic tensor decomposition JMLR:v15:anandkumar14b . For brevity, we defer the algorithmic details to Appendix A.
Next, we give a rigorous definition of the transformer model. To maintain consistency with existing literature, we adopt the notational conventions presented in bai2023tfstats , with modifications tailored to our specific context. We consider a sequence of $N$ input vectors $\{h_{i}\}_{i=1}^{N}\subset\mathbb{R}^{D}$, which can be compactly represented as an input matrix ${\mathbf{H}}=[h_{1},\dots,h_{N}]\in\mathbb{R}^{D\times N}$, where each $h_{i}$ corresponds to a column of ${\mathbf{H}}$ (also referred to as a token).
First, we introduce the definition of the attention layer.
Definition 1 (Attention layer).
A (self-)attention layer with $M$ heads is denoted as ${\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}}(\cdot)$ with parameters ${\boldsymbol{\Theta}}_{\tt attn}={\{({\mathbf{V}}_{m},{\mathbf{Q}}_{m},{\mathbf{K}}_{m})\}}_{m\in[M]}\subset\mathbb{R}^{D\times D}$. On any input sequence ${\mathbf{H}}\in\mathbb{R}^{D\times N}$,
$\widetilde{{\mathbf{H}}}={\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}}({\mathbf{H}})\mathrel{\mathop{:}}={\mathbf{H}}+\sum_{m=1}^{M}({\mathbf{V}}_{m}{\mathbf{H}})\text{SoftMax}{\left(({\mathbf{K}}_{m}{\mathbf{H}})^{\top}({\mathbf{Q}}_{m}{\mathbf{H}})\right)}\in\mathbb{R}^{D\times N},$
In vector form,
$\widetilde{{\mathbf{h}}}_{i}={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}}({\mathbf{H}})\right]}_{i}={\mathbf{h}}_{i}+\sum_{m=1}^{M}\sum_{j=1}^{N}{\left[\text{SoftMax}{\left({\left({\left({\mathbf{Q}}_{m}{\mathbf{h}}_{i}\right)}^{\top}{\left({\mathbf{K}}_{m}{\mathbf{h}}_{j}\right)}\right)}_{j=1}^{N}\right)}\right]}_{j}{\mathbf{V}}_{m}{\mathbf{h}}_{j}.$
Here SoftMax is the activation function defined by ${\text{SoftMax}}{\left(v\right)}={\left(\frac{\exp(v_{1})}{\sum_{i=1}^{d}\exp(v_{i})},\cdots,\frac{\exp(v_{d})}{\sum_{i=1}^{d}\exp(v_{i})}\right)}$ for $v\in\mathbb{R}^{d}$.
The Multilayer Perceptron(MLP) layer is defined as follows.
Definition 2 (MLP layer).
A (token-wise) MLP layer with hidden dimension $D^{\prime}$ is denoted as $\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}}(\cdot)$ with parameters ${\boldsymbol{\Theta}}_{\tt mlp}=({\mathbf{W}}_{1},{\mathbf{W}}_{2})\in\mathbb{R}^{D^{\prime}\times D}\times\mathbb{R}^{D\times D^{\prime}}$. On any input sequence ${\mathbf{H}}\in\mathbb{R}^{D\times N}$,
$\widetilde{{\mathbf{H}}}=\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}}({\mathbf{H}})\mathrel{\mathop{:}}={\mathbf{H}}+{\mathbf{W}}_{2}\sigma({\mathbf{W}}_{1}{\mathbf{H}}),$
where $\sigma:\mathbb{R}\to\mathbb{R}$ is the ReLU function. In vector form, we have $\widetilde{{\mathbf{h}}}_{i}={\mathbf{h}}_{i}+{\mathbf{W}}_{2}\sigma({\mathbf{W}}_{1}{\mathbf{h}}_{i})$.
Then we can use the above definitions to define the transformer model.
Definition 3 (Transformer).
An $L$-layer transformer, denoted as ${\rm TF}_{{\boldsymbol{\Theta}}_{\tt TF}}(\cdot)$, is a composition of $L$ self-attention layers each followed by an MLP layer:
${\rm TF}_{{\boldsymbol{\Theta}}_{\tt TF}}({\mathbf{H}})=\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(L)}}{\left({\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(L)}}{\left(\cdots\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(1)}}{\left({\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(1)}}{\left({\mathbf{H}}\right)}\right)}\right)}\right)}.$
Here the parameter ${{\boldsymbol{\Theta}}_{\tt TF}}=({\boldsymbol{\Theta}}_{\tt attn}^{(1:L)},{\boldsymbol{\Theta}}_{\tt mlp}^{(1:L)})$ consists of the attention layers ${\boldsymbol{\Theta}}_{\tt attn}^{(\ell)}={\{({\mathbf{V}}^{(\ell)}_{m},{\mathbf{Q}}^{(\ell)}_{m},{\mathbf{K}}^{(\ell)}_{m})\}}_{m\in[M^{(\ell)}]}\subset\mathbb{R}^{D\times D}$, the MLP layers ${\boldsymbol{\Theta}}_{\tt mlp}^{(\ell)}=({\mathbf{W}}^{(\ell)}_{1},{\mathbf{W}}^{(\ell)}_{2})\in\mathbb{R}^{D^{(\ell)}\times D}\times\mathbb{R}^{D\times D^{(\ell)}}$.
2.2 The TGMM architecture
A recent line of work xie2021explanation ; garg2022can ; bai2023tfstats ; akyrek2023what ; li2023transformers has been studying the capability of transformer that functions as a data-driven algorithm under the context of in-context learning (ICL). However, in contrast to the setups therein where inputs consist of both features and labels, under the unsupervised GMM setup, there is no explicitly provided label information. Therefore, we formulate the learning problem as learning an estimation algorithm instead of learning a prediction algorithm as in the case of ICL. A notable property of GMM is that the structure of the estimand depends on an unknown parameter $K$, which is often treated as a hyper-parameter in GMM estimation titterington_85 ; mclachlan2000finite . For clarity of representation, we define an isotropic Gaussian mixture task as $\mathcal{T}=({\boldsymbol{\theta}},{\mathbf{X}},K)$, where ${\mathbf{X}}$ is a i.i.d. sample generated according to ground truth ${\boldsymbol{\theta}}$ according to the isotropic GMM law and $K$ is the configuration used during estimation which we assume to be the same as the number of components of the ground truth ${\boldsymbol{\theta}}$. The GMM task is solved via applying some algorithm $\mathcal{A}$ that takes ${\mathbf{X}}$ and $K$ as inputs and outputs an estimate of the ground truth $\widehat{{\boldsymbol{\theta}}}=\mathcal{A}({\mathbf{X}};K)$.
In this paper, we propose a transformer-based architecture, transformers-for-Gaussian-mixtures (TGMM), as a GMM task solver that allows flexibility in its outputs, while at the same time being parameter-efficient, as illustrated in Fig. 1: A TGMM model supports solving $s$ different GMM tasks with $K\in\mathcal{K}:={\{K_{1},\ldots,K_{s}\}}$. Given inputs $N$ data points ${\mathbf{X}}$ and a structure configuration of the estimand $K$. TGMM first augments the inputs with auxiliary configurations about $K$ via concatenating it with a task embedding ${\mathbf{P}}=\text{embed}(K)$, i.e., ${\mathbf{H}}=\left[{\mathbf{X}}||{\mathbf{P}}\right]$, and use a linear $\operatorname{Readin}$ layer to project project the augmented inputs onto a shared hidden representation space for several estimand structures ${\{K_{1},\ldots,K_{s}\}}$, which is then manipulated by a shared transformer backbone that produces task-aware hidden representations. The TGMM estimates are then decoded by task-specific $\operatorname{Readout}$ modules. More precisely, with target decoding parameters of $K$ components, the $\operatorname{Readout}$ module first perform an attentive-pooling operation lee2019set :
$\mathbf{O}=({\mathbf{V}}_{o}{\mathbf{H}})\text{SoftMax}{\left(({\mathbf{K}}_{o}{\mathbf{H}})^{\top}{\mathbf{Q}}_{o}\right)}\in\mathbb{R}^{(d+K)\times K},$
where ${\mathbf{V}}_{o},{\mathbf{K}}_{o}\in\mathbb{R}^{(d+K)\times D}$, ${\mathbf{Q}}_{o}\in\mathbb{R}^{(d+K)\times K}$. The estimates for mixture probability are then extracted by a row-wise mean-pooling of the first $K$ rows of $\mathbf{O}$, and the estimates for mean vectors are the last $d$ rows of $\mathbf{O}$. We wrap the above procedure as ${\{\widehat{\pi}_{k},\widehat{\mu}_{k}\}}_{i\in[K]}=\operatorname{Readout}_{{\boldsymbol{\Theta}}_{\tt out}}{\left({\mathbf{H}}\right)}$. TGMM is parameter-efficient in the sense that it only introduces extra parameter complexities of the order $O(sdD)$ in addition to the backbone. We give a more detailed explanation of the parameter efficiency of TGMM in appendix Appendix B. We wrap the TGMM model into the following form:
${\rm TGMM}_{\boldsymbol{\Theta}}({\mathbf{X}};K)=\operatorname{Readout}_{{\boldsymbol{\Theta}}_{\tt out}}{\left({\rm TF}_{{\boldsymbol{\Theta}}_{\tt TF}}{\left(\operatorname{Readin}_{{\boldsymbol{\Theta}}_{\tt in}}{\left(\left[{\mathbf{X}}||\text{embed}(K)\right]\right)}\right)}\right)}.$
Above, the parameter ${\boldsymbol{\Theta}}=({{\boldsymbol{\Theta}}_{\tt TF}},{\boldsymbol{\Theta}}_{\tt in},{\boldsymbol{\Theta}}_{\tt out})$ consists of the parameters in the transformer ${{\boldsymbol{\Theta}}_{\tt TF}}$ and the parameters in the $\operatorname{Readin}$ and the $\operatorname{Readout}$ functions ${\boldsymbol{\Theta}}_{\tt in}$, ${\boldsymbol{\Theta}}_{\tt out}$.
Figure 1: Illustration of the proposed TGMM architecture: TGMM utilizes a shared transformer backbone that supports solving $s$ different kind of GMM tasks via a task-specific $\operatorname{Readout}$ strategies.
2.3 Meta training procedure
We adopt the meta-training framework as in garg2022can ; bai2023tfstats and utilize diverse synthetic tasks to learn the TGMM model. In particular, during each step of the learning process, we first use a TaskSampler routine (described in Algorithm 1) to generate a batch of $n$ tasks, with each task having a probably distinct sample size. The TGMM model outputs estimates for each task, i.e., ${\{\widehat{\mu}_{k},\widehat{\pi}_{k}\}}_{k\in[K]}={\rm TGMM}_{{\boldsymbol{\Theta}}}{\left({\mathbf{X}};K\right)}$. Define $\widehat{{\boldsymbol{\pi}}}\mathrel{\mathop{:}}={\{\widehat{\pi}_{k}\}}_{k\in[K]}$ and $\widehat{{\boldsymbol{\mu}}}\mathrel{\mathop{:}}={\{\widehat{\mu}_{k}\}}_{k\in[K]}$. For a batch of tasks ${\{\mathcal{T}_{i}\}}_{i\in[n]}={\{{\mathbf{X}}_{i},{\boldsymbol{\theta}}_{i},K_{i}\}}_{i\in[n]}$, denote by ${\boldsymbol{\theta}}_{i}={\boldsymbol{\mu}}_{i}\cup{\boldsymbol{\pi}}_{i}$ and $\widehat{{\boldsymbol{\theta}}}_{i}=\widehat{{\boldsymbol{\mu}}}_{i}\cup\widehat{{\boldsymbol{\pi}}}_{i}={\rm TGMM}_{{\boldsymbol{\Theta}}}{\left({\mathbf{X}}_{i};K_{i}\right)}$, $i\in[n]$. Then the learning objective is thus:
$\widehat{L}_{n}{\left({\boldsymbol{\Theta}}\right)}=\frac{1}{n}\sum_{i=1}^{n}\ell_{\mu}(\widehat{{\boldsymbol{\mu}}}_{i},{\boldsymbol{\mu}}_{i})+\ell_{\pi}(\widehat{{\boldsymbol{\pi}}}_{i},{\boldsymbol{\pi}}_{i}).$ (2)
where $\ell_{\mu}$ and $\ell_{\pi}$ are loss functions for estimation of $\mu$ and $\pi$, respectively. We will by default use square loss for $\ell_{\mu}$ and cross entropy loss for $\ell_{\pi}$. Note that the task sampling procedure relies on several sampling distributions $p_{\mu},p_{\pi},p_{N},p_{K}$, which are themselves dependent upon some global configurations such as the dimension $d$ as well as the task types $\mathcal{K}$. We will omit those dependencies on global configurations when they are clear from context. The (meta) training procedure is detailed in Algorithm 2.
Algorithm 1 TaskSampler
1:sampling distributions $p_{\mu},p_{\pi},p_{N},p_{K}$.
2:Sample the type of task (i.e., number of mixture components) $K\sim p_{K}$.
3:Sample a GMM task according to the type of task
[…]
4:Sample the size of inputs $N\sim p_{N}$.
[…]
6:return An (isotropic) GMM task $\mathcal{T}=({\mathbf{X}},{\boldsymbol{\theta}},K)$.
Algorithm 2 (Meta) Training procedure for TGMM
1:task dimension $d$, task types $\mathcal{K}=\{K_{1},\ldots,K_{s}\}$, number of tasks per step $n$, number of steps $T$.
2:Initialize a TGMM model ${\rm TGMM}_{{\boldsymbol{\Theta}}^{(0)}}$.
3:for $t=1:T$ do
4:     Sample $n$ tasks ${\{\mathcal{T}_{i}\}}_{i\in[n]}$ independently using the TaskSampler routine 1.
5:     Compute the training objective $\widehat{L}_{n}{\left({\boldsymbol{\Theta}}^{(t-1)}\right)}$ as in (2).
6:     Update ${\boldsymbol{\Theta}}^{(t-1)}$ into ${\boldsymbol{\Theta}}^{(t)}$ using any gradient based training algorithm like AdamW.
7:end for
8:return Trained model ${\rm TGMM}_{{\boldsymbol{\Theta}}^{(T)}}$.
3 Experiments
In this section, we empirically investigate TGMM’s capability of learning to solve GMMs. We focus on the following research questions (RQ):
RQ1 Effectiveness: How well do TGMM solve GMM problems, compared to classical algorithms?
RQ2 Robustness: How well does TGMM perform over test tasks unseen during training?
RQ3 Flexibility: Can we extend the current formulation by adopting alternative backbone architectures or relaxing the isotropic setting to more sophisticated models like anisotropic GMM?
3.1 Experimental Setup
We pick the default backbone of TGMM similar to that in garg2022can ; bai2023tfstats , with a GPT-2 type transformer encoder radford2019language of $12$ layers, $4$ heads, and $128$-dimensional hidden state size. The task embedding dimension is fixed at $128$. Across all the experiments, we use AdamW loshchilov2017fixing as the optimizer and use both learning rate and weight decay coefficient set to $10^{-4}$ without further tuning. During each meta-training step, we fix the batch size to be $64$ and train $10^{6}$ steps. For the construction of TaskSampler, the sampling distributions are defined as follows: For $p_{K}$, We sample $K$ uniformly from ${\{2,3,4,5\}}$; For $p_{\mu}$, given dimension $d$ and number of components $K$, we sample each component uniformly from $[-5,5]^{d}$. Additionally, to prevent collapsed component means ndaoud2022sharp , we filter the generated mean vectors with a maximum pairwise cosine similarity threshold of $0.8$; For $p_{\pi}$, given $K$, we sample each $\pi_{k}$ uniformly from $[0.2,0.8]$ and normalize them to be a probability vector; For $p_{N}$, Given a maximum sample size $N_{0}$, we sample $N$ uniformly from $[N_{0}/2,N_{0}]$. The default choice of $N_{0}$ is $128$. During evaluation, we separately evaluate $4$ tasks with $2,3,4,5$ components, respectively. With a sample size of $128$ and averaging over $1280$ randomly sampled tasks.
Metrics. We use $\ell_{2}$-error as evaluation metrics in the experiments. We denote the output of the TGMM as $\widehat{{\boldsymbol{\theta}}}\mathrel{\mathop{:}}={\{\widehat{\pi}_{1},\widehat{\mu}_{1},\widehat{\pi}_{2},\widehat{\mu}_{2},\cdots\widehat{\pi}_{K},\widehat{\mu}_{K}\}}$. The rigorous definition is
$\frac{1}{K}\sum_{k\in[K]}{\left(\frac{1}{d}\left\|{\widehat{\mu}_{\tilde{\sigma}(i)}-\mu_{i}}\right\|^{2}+{\left(\widehat{\pi}_{\tilde{\sigma}(i)}-\pi_{i}\right)}^{2}\right)},$
where $\tilde{\sigma}$ is the permutation such that $\tilde{\sigma}=\arg\min_{\sigma}\sum_{k\in[K]}\left\|{\widehat{\mu}_{{\sigma}(i)}-\mu_{i}}\right\|^{2}$. We obtain the permutation via solving a linear assignment program using the Jonker-Volgenant algorithm crouse2016implementing . We also report all the experimental results under two alternative metrics: cluster-classification accuracy and log-likelihood in Section E.2.
3.2 Results and findings
RQ1: Effectiveness
We compare the performance of a learned TGMM with the classical EM algorithm and spectral algorithm under $4$ scenarios where the problem dimension ranges over ${\{2,8,32,128\}}$. The results are reported in Fig. 2. We observe that all three algorithms perform competitively (reaching almost zero estimation error) when $K=2$. However, as the estimation problem gets more challenging as $K$ increases, the EM algorithm gets trapped in local minima and underperforms both spectral and TGMM. Moreover, while the spectral algorithm performs comparably with TGMM, it cannot handle cases when $K>d$, which is effectively mitigated by TGMM, with corresponding performances surpassing those of the EM algorithm. This demonstrates the effectiveness of TGMM for learning an estimation algorithm that efficiently solves GMM problems.
Figure 2: Performance comparison between TGMM and two classical algorithms, reported in $\ell_{2}$-error.
RQ2: Robustness To assess the robustness of the learned TGMM, we consider two types of test-time distribution shifts:
Shifts in sample size $N$ Under this scenario, we evaluate the learned TGMM model on tasks with sample size $N^{\text{test}}$ that are unseen during training.
Shifts in sampling distributions Under this scenario, we test the learned TGMM model on tasks that are sampled from different sampling distributions that are used during training. Specifically, we use the same training sampling configuration as stated in Section 3.1 and test on the following perturbed sampling scheme, with $\tilde{\mu}_{k}=\mu_{k}+\sigma_{p}\varepsilon_{k}$, where $\mu_{k}\overset{i.i.d.}{\sim}\operatorname{Unif}{\left([-5,5]^{d}\right)}$, $\varepsilon_{k}\overset{i.i.d.}{\sim}{\mathcal{N}}{\left(0,I_{d}\right)}$, $k\in[K]$ and ${\{\varepsilon_{k}\}}_{k\in[K]}$ is independent with ${\{\mu_{k}\}}_{k\in[K]}$.
Figure 3: Assessments of TGMM under test-time task distribution shifts I: A line with $N_{0}^{\text{train}}\rightarrow N^{\text{test}}$ draws the performance of a TGMM model trained over tasks with sample size randomly sampled in $[N_{0}^{\text{train}}/2,N_{0}^{\text{train}}]$ and evaluated over tasks with sample size $N^{\text{test}}$. We can view the configuration $128\rightarrow 128$ as an in-distribution test and the rest as out-of-distribution tests.
In Fig. 3, we report the assessments regarding shifts in sample size, where we set $N_{\text{test}}$ to be $128$ and vary the training configuration $N_{0}$ to range over ${\left\{32,64,128\right\}}$, respectively. The results demonstrate graceful performance degradation of out-of-domain testing performance in comparison to the in-domain performance. To measure performance over shifted test-time sampling distributions, we vary the perturbation scale $\sigma_{p}\in\{0,1,\ldots,10\}$ with problem dimension fixed at $d=8$. The results are illustrated in Fig. 4 along with comparisons to EM and spectral baselines. As shown in the results, with the increase of the perturbation scale, the estimation problem gets much harder. Nevertheless, the learned TGMM can still outperform the EM algorithm when $K>2$. Both pieces of evidence suggest that our meta-training procedure indeed learns an algorithm instead of overfitting to some training distribution.
Figure 4: Assessments of TGMM under test-time task distribution shifts II: $\ell_{2}$-error of estimation when the test-time tasks $\mathcal{T}^{\text{test}}$ are sampled using a mean vector sampling distribution $p_{\mu}^{\text{test}}$ different from the one used during training.
RQ3: Flexibility Finally, we initiate two studies that extend both the TGMM framework and the (meta) learning problem of solving isotropic GMMs. In our first study, we investigated alternative architectures for the TGMM backbone. Motivated by previous studies park2024can that demonstrate the in-context learning capability of linear attention models such as Mamba series gu2023mamba ; dao2024transformers . We test replacing the backbone of TGMM with a Mamba2 dao2024transformers model with its detailed specifications and experimental setups listed in Section E.1. The results are reported in Fig. 5, suggesting that while utilizing Mamba2 as the TGMM backbone still yields non-trivial estimation efficacy, it is in general inferior to transformer backbone under comparable model complexity.
In our second study, we adapted TGMM to be compatible with more sophisticated GMM tasks via relaxing the isotropic assumption. Specifically, we construct anisotropic GMM tasks via equipping it with another scale sampling mechanism $p_{\sigma}$, where for each task we sample $\sigma\sim\text{softplus}(\tilde{\sigma})$ with $\tilde{\sigma}$ being sampled uniformly from $[-1,1]^{d}$. We adjust the output structure of TGMM accordingly so that its outputs can be decoded into both estimates of both mean vectors, mixture probabilities, and scales, which are detailed in Section E.1. Note that the spectral algorithm does not directly apply to anisotropic setups, limiting its flexibility. Consequently, we compare TGMM with the EM approach and plot results in Fig. 6 with the $\ell_{2}$-error metric accommodating errors from scale estimation. The results demonstrate a similar trend as in evaluations in the isotropic case, showcasing TGMM as a versatile tool in GMM learning problems.
Additional experiments We postpone some further evaluations to Appendix E, where we present a complete report consisting of more metrics and conduct several ablations on the effects of backbone scales and sample sizes.
Figure 5: Performance comparisons between TGMM using transformer and Mamba2 as backbone, reported in $\ell_{2}$-error.
Figure 6: Performance comparison between TGMM and the EM algorithm on anisotropic GMM tasks, reported in $\ell_{2}$-error.
4 Theoretical understandings
In this section, we provide some theoretical understandings for the experiments.
4.1 Understanding TGMM
We investigate the expressive power of transformers-for-Gaussian-mixtures(TGMM) as demonstrated in Section 3. Our analysis presents two key findings that elucidate the transformer’s effectiveness for GMM estimation: 1. Transformer can approximate the EM algorithm; 2. Transformer can approximate the power iteration of cubic tensor.
Transformer can approximate the EM algorithm. We show that transformer can efficiently approximate the EM algorithm (Algorithm A.1; see Appendix A) and estimate the parameters of GMM. Moreover, we show that transformer with one backbone can handle tasks with different dimensions and components simultaneously. The formal statement appears in Appendix C due to space limitations.
Theorem 1 (Informal).
There exists a $2L$-layer transformer ${\rm TF}_{\boldsymbol{\Theta}}$ such that for any $d\leq d_{0}$, $K\leq K_{0}$ and task ${\mathcal{T}}={\left({\mathbf{X}},{\boldsymbol{\theta}},K\right)}$ satisfying some regular conditions, given suitable embeddings, ${\rm TF}_{\boldsymbol{\Theta}}$ approximates EM algorithm $L$ steps and estimates ${\boldsymbol{\theta}}$ efficiently.
Transformer can approximate power iteration of cubic tensor. Since directly implementing the spectral algorithm with transformers proves prohibitively complex, we instead demonstrate that transformers can effectively approximate its core computational step–the power iteration for cubic tensors (Algorithm 1 in JMLR:v15:anandkumar14b ; see Appendix A). Specifically, we prove that a single-layer transformer can approximate the iteration step:
$v^{(j+1)}=T{\left(I,v^{(j)},v^{(j)}\right)},~{}j\in\mathbb{N},$ (4)
where $I$ denotes the identity matrix and $T$ represents the given cubic tensor. For technical tractability, we assume the attention layer employs a ReLU activation function. The formal statement appears in Appendix D due to space limitations.
Theorem 2 (Informal).
There exists a $2L$-layer transformer ${\rm TF}_{\boldsymbol{\Theta}}$ with ReLU activation such that for any $d\leq d_{0}$, $T\in\mathbb{R}^{d\times d\times d}$ and $v^{(0)}\in\mathbb{R}^{d}$, given suitable embeddings, ${\rm TF}_{\boldsymbol{\Theta}}$ implements $L$ steps of Eq. 4 exactly.
We give some discussion of the theorems in the following remarks.
Remark 1.
(1) Theorem 1 demonstrates that a transformer architecture can approximate the EM algorithm for GMM tasks with varying numbers of components using a single shared set of parameters (i.e., one backbone ${\boldsymbol{\Theta}}$). This finding supports the empirical effectiveness of TGMM (RQ1 in Section 3.2). Additionally, Theorem 2 establishes that transformers can approximate power iterations for third-order tensors across different dimensions, further corroborating the model’s ability to generalize across GMMs with varying component counts.
(2) Theorem 1 holds uniformly over sample sizes $N$ and sampling distributions under mild regularity conditions, aligning with the observed robustness of TGMM (RQ2 in Section 3.2).
Remark 2.
Different "readout" functions are also required to extract task-specific parameters in our theoretical analysis, aligning with the architectural design described in Section 2.2. For further discussion, refer to Remark C.3 in Section C.2.
4.2 Proof Ideas
Proof Idea of Theorem 1. We present a brief overview of the proof strategy for Theorem 1. Our approach combines three key components: (1) the convergence properties of the population-EM algorithm pmlr-v125-kwon20a , (2) concentration bounds between population and sample quantities (established via classical empirical process theory), and (3) a novel transformer architecture construction. The transformer design is specifically motivated by the weighting properties of the softmax activation function, which naturally aligns with the EM algorithm’s update structure. For intuitive understanding, Fig. 7 provides a graphical illustration of this construction. The full proof is in Appendix C.
Figure 7: (Informal version)Transformer Construction for Approximating EM Algorithm Iterations. The word "clean" means setting all positions of the corresponding vector to zero.
Proof Idea of Theorem 2. To approximate Eq. 4, we perform a two-dimensional computation within a single-layer transformer. The key idea is to leverage the number of attention heads $M$ to handle one dimension while utilizing the $Q,K,V$ structure in the attention layer. Specifically, let $T=(T_{i,j,m})_{i,j,m\in[d]}$ and $v^{(j)}=(v_{i}^{(j)})_{i\in[d]}$. Then, Eq. 4 can be rewritten as
$v^{(j+1)}=\sum_{j,m\in[d]}v_{j}v_{l}T_{:,j,m},$
where $T_{:,j,m}=(T_{i,j,m})_{i\in[d]}\in\mathbb{R}^{d}$. This operation can be implemented using $d$ attention heads, where each head processes a dimension of size $d$ (Fig. 8). The complete construction and proof are provided in Appendix D.
Figure 8: Illustration of implementing Eq. 4 via a multi-head attention structure, where colored boxes denote corresponding implementation components. Here $\sigma$ denotes the ReLU function.
5 Conclusion and discussions
In this paper, we investigate the capabilities of transformers in GMM tasks from both theoretical and empirical perspectives. We discuss the limitations and potential future research directions here.
Limitations and future work directions. First, while our theoretical analysis focuses on the approximation ability of transformers, the optimization dynamics remain unexplored. Second, approximating the full spectral algorithm (Algorithm A.2; see Appendix A) presents a significant challenge, which we leave for future work. Third, our study is limited to the expressivity of transformers on classical GMM tasks; exploring their performance on other unsupervised learning tasks is an interesting direction that warrants further investigation.
Appendix
[…]
Organization of the Appendix. In Appendix A, we formally present the GMM algorithms referenced in Section 2. We discuss the parameter efficiency of TGMM in Appendix B. Rigorous statements and proofs of Theorem 1 and Theorem 2 are provided in Appendix C and Appendix D, respectively. Additional experimental details are included in Appendix E.
Additional notations in the Appendix. The maximum between two scalars $a,b$ is denoted as $a\vee b$. For a vector $v\in\mathbb{R}^{d}$, let $\left\|{v}\right\|_{\infty}\mathrel{\mathop{:}}=\max_{i\in[d]}|v_{i}|$ be its infinity norm. We use ${\mathbf{0}}_{d}$ to denote the zero vector and ${\mathbf{e}}_{i}\in\mathbb{R}^{d}$ to denote the $i$-th standard unit vector in $\mathbb{R}^{d}$. For a matrix ${\mathbf{A}}\in\mathbb{R}^{d_{1}\times d_{2}}$, we denote $\left\|{{\mathbf{A}}}\right\|_{2}\mathrel{\mathop{:}}=\sup_{\left\|{x}\right\|_{2}=1}\left\|{{\mathbf{A}}x}\right\|$ as its operator norm. We use $\widetilde{O}{\left(\cdot\right)}$ to denote $O{\left(\cdot\right)}$ with hidden $\log$ factors. For clarify, we denote the ground-truth parameters of GMM with a superscript , i.e. ${\{\pi_{k}^{\ast},\mu_{k}^{\ast}\}}_{k\in[K]}$, throughout this appendix.
Appendix A Algorithm Details
We state the classical algorithms of GMM mention in Section 2 in this section.
Algorithm A.1 EM algorithm for GMM
1:$\{X_{i},i\in[N]\}$, ${\boldsymbol{\theta}}^{(0)}=\{\pi_{1}^{(0)},\mu_{1}^{(0)},\cdots\pi_{K}^{(0)},\mu_{K}^{(0)}\}$
2:$j\leftarrow 0$
3:while not converge do
4:     E-step: $w_{k}^{(j+1)}(X_{i})=\frac{\pi_{k}^{(j)}\phi(X_{i};\mu_{k}^{(j)})}{\sum_{k\in[K]}\pi_{k}^{(j)}\phi(X_{i};\mu_{k}^{(j)})}$, $i\in[N]$, $k\in[K]$
5:     M-step: $\pi_{k}^{(j+1)}=\frac{\sum_{i\in[N]}w_{k}^{(j+1)}(X_{i})}{N}$, $\mu_{k}^{(j+1)}=\frac{\sum_{i\in[N]}w_{k}^{(j+1)}(X_{i})X_{i}}{\sum_{i\in[N]}w_{k}^{(j+1)}(X_{i})}$, $k\in[K]$
6:     $j\leftarrow j+1$
7:end while
Algorithm A.2 Spectral Algorithm for GMM
1:$\{X_{i},i\in[N]\}$
2:Compute the empirical moments $\hat{M}_{2}$ and $\hat{M}_{3}$ by
$\hat{M}_{2}$ $=\frac{1}{N}\sum_{i\in[N]}X_{i}\otimes X_{i}-I_{d},~{}$
$\hat{M}_{3}$ $=\frac{1}{N}\sum_{i\in[N]}X_{i}\otimes X_{i}\otimes X_{i}-\frac{1}{N}\sum_{i\in[N],j\in[d]}{\left(X_{i}\otimes{\mathbf{e}}_{j}\otimes{\mathbf{e}}_{j}+{\mathbf{e}}_{j}\otimes X_{i}\otimes{\mathbf{e}}_{j}+{\mathbf{e}}_{j}\otimes{\mathbf{e}}_{j}\otimes X_{i}\right)}$
3:Do first $K$-th singular value decomposition(SVD) for $\hat{M}_{2}$: $\hat{M}_{2}\approx UDU^{\top}$ and let $W=UD^{-1/2}$, $B=UD^{1/2}$
4:Do first $K$-th robust tensor decomposition (Algorithm 1 in [3], see Algorithm A.3) for $\tilde{M}_{3}=\hat{M}_{3}{\left(W,W,W\right)}$:
$\tilde{M}_{3}\approx\sum_{k\in[K]}\lambda_{k}v_{k}^{\otimes 3}$
return $\hat{\pi}_{k}=\lambda_{k}^{-2}$, $\hat{\mu}_{k}=\lambda_{k}Bv_{k}$, $k\in[K]$.
Algorithm A.3 Robust Tensor Power Method
1:symmetric tensor $T\in\mathbb{R}^{d\times d\times d}$, number of iterations $L$, $N$.
2:the estimated eigenvector/eigenvalue pair; the deflated tensor.
3:for $\tau=1$ to $L$ do
4:     Draw $v_{0}^{(\tau)}$ uniformly at random from the unit sphere in $\mathbb{R}^{d}$.
5:     for $t=1$ to $N$ do
6:         Compute power iteration update:
7:         $v_{t}^{(\tau)}:=\frac{T(I,v_{t-1}^{(\tau)},v_{t-1}^{(\tau)})}{\|T(I,v_{t-1}^{(\tau)},v_{t-1}^{(\tau)})\|}$
8:     end for
9:end for
10:Let $\tau^{*}:=\arg\max_{\tau\in[L]}\{T(v_{N}^{(\tau)},v_{N}^{(\tau)},v_{N}^{(\tau)})\}$.
11:Do $N$ power iteration updates (line 5) starting from $v_{N}^{(\tau^{*})}$ to obtain $\hat{v}$.
12:Set $\hat{\lambda}:=\tilde{T}(\hat{v},\hat{v},\hat{v})$.
13:return the estimated eigenvector/eigenvalue pair $(\hat{v},\hat{\lambda})$; the deflated tensor $\tilde{T}-\hat{\lambda}\hat{v}^{\otimes 3}$.
Appendix B On the parameter efficiency of TGMM
Aside from its backbone, the extra parameters in a TGMM comprises the following:
Parameters in the task embedding module
This part has a parameter count of $s\times d_{\text{task}}$.
Parameters in the $\operatorname{Readin}$ layer
This part has a parameter count of $O((d_{\text{task}}+d)\times D)$.
Parameters in the $\operatorname{Readout}$ layer
This part has a parameter count of $O(sdD)$, which comprises of parameters from $s$ distinct attention mechanisms.
As $d_{\text{task}}$ is typically of the order $O(D)$, we conclude that the total extra parameter complexity is of the order $O(sdD)$, which in practice is often way smaller than the parameter complexity of the backbone, i.e., of the order $O(LD^{2})$ Meanwhile, a naive implementation of adapting transformer architecture to solve $s$ distinct GMM tasks require a different transformer backbone. As the complexity of backbone often dominate those of extra components, the TGMM implementation can reduce the parameter complexity by an (approximate) factor of $1/s$ in practice.
Appendix C Formal statement of Theorem 1 and proofs
For analytical tractability, we implement $\operatorname{Readin}$ as an identity transformation and define $\operatorname{Readout}$ to extract targeted matrix elements hence they are both fixed functions. Actually, we also need "Readout" functions to get the estimated parameters for different tasks, see Remark C.3. To theoretical convenience, we use the following norm of transformers, which differs slightly from the definition in [4].
${\left|\!\left|\!\left|{{\boldsymbol{\Theta}}}\right|\!\right|\!\right|}\mathrel{\mathop{:}}=\max_{\ell\in[L]}{\left\{\max_{m\in[M]}{\left\{\left\|{{\mathbf{Q}}_{m}^{(\ell)}}\right\|_{2},\left\|{{\mathbf{K}}_{m}^{(\ell)}}\right\|_{2},\left\|{{\mathbf{V}}_{m}^{(\ell)}}\right\|_{2}\right\}}+\left\|{{\mathbf{W}}_{1}^{(\ell)}}\right\|_{2}+\left\|{{\mathbf{W}}_{2}^{(\ell)}}\right\|_{2}\right\}}.$
Then the transformer class can be defined as
${\mathcal{F}}\mathrel{\mathop{:}}={\mathcal{F}}(L,D,D^{\prime},M,B_{\boldsymbol{\Theta}})={\left\{{\rm TF}_{\boldsymbol{\Theta}},{\left|\!\left|\!\left|{{\boldsymbol{\Theta}}}\right|\!\right|\!\right|}\leq B_{\boldsymbol{\Theta}},D^{(\ell)}\leq D^{\prime},M^{(\ell)}\leq M,\ell\in[L]\right\}}.$
C.1 Formal statement of Theorem 1
First, we introduce some notations. We define $\pi_{\min}=\min_{i}\pi_{i}^{\ast}$, $\rho_{\pi}=\max_{i}\pi_{i}^{*}/\min_{i}\pi_{i}^{*}$. We use $R_{ij}=\|\mu_{i}^{\ast}-\mu_{j}^{\ast}\|$ to denote the pairwise distance between components and $R_{\min}=\min_{i\neq j}R_{ij}$, $R_{max}={\left(\max_{i\neq j}R_{ij}\right)}\vee{\left(\max_{i\in[K]}\left\|{\mu_{i}^{\ast}}\right\|\right)}$. Without the loss of generality, we assume that $R_{\max}\geq 1$. For dimension and components adaptation, we assume $d\leq d_{0}$ and $K\leq K_{0}$. Since in practice the sample size $N$ is much larger than the number of components $K$, we assume that $N$ is divisible by $K$, i.e. $N/K\in\mathbb{N}$. Otherwise, we only consider the first $K\lfloor N/K\rfloor$ samples and drop the others. We encode ${\mathbf{X}}=\{X_{i}\}_{i=1}^{N}$ into an input sequence ${\mathbf{H}}$ as the following:
${\mathbf{H}}=\begin{bmatrix}\overline{X}_{1}&\overline{X}_{2}&\dots&\overline{X}_{N}\\ {\mathbf{p}}_{1}&{\mathbf{p}}_{2}&\dots&{\mathbf{p}}_{N}\end{bmatrix}\in\mathbb{R}^{D\times N},~{}{\mathbf{p}}_{i}=\begin{bmatrix}\overline{{\boldsymbol{\theta}}}_{i}\\ {\bf r}_{i}\end{bmatrix},~{}\overline{{\boldsymbol{\theta}}}_{i}=\begin{bmatrix}\overline{{\boldsymbol{\pi}}}_{\log}\\ \overline{\mu}_{i\%K}\\ c_{i\%K}\\ {\mathbf{0}}_{3K_{0}}\end{bmatrix}\in\mathbb{R}^{d_{0}+4K_{0}+1},~{}{\bf r}_{i}=\begin{bmatrix}{\mathbf{0}}_{\tilde{D}}\\ 1\\ {\mathbf{e}}_{i\%K}\end{bmatrix}\in\mathbb{R}^{D-(2d_{0}+3K_{0}+1)},$ (5)
where $\overline{X}_{i}=[{X}_{i}^{\top},{\mathbf{0}}_{d_{0}-d}^{\top}]^{\top}$, $\overline{{\boldsymbol{\pi}}}_{\log}=[{\boldsymbol{\pi}}_{\log}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $\overline{\mu}_{i\%K}=[\mu_{i\%K}^{\top}$, ${\mathbf{0}}_{d_{0}-d}^{\top}]^{\top}$, $c_{i\%K}\in\mathbb{R}$ and ${\mathbf{e}}_{i\%K}\in\mathbb{R}^{K_{0}}$ denotes the ${i\%K}$-th standard unit vector. To match the dimension, $\tilde{D}=D-(2d_{0}+5K_{0}+2)$. We choose $D=O(d_{0}+K_{0})$ to get the encoding above. For the initialization, we choose ${\boldsymbol{\pi}}_{\log}=\log{\boldsymbol{\pi}}^{(0)}$, $\mu_{i}=\mu_{i}^{(0)}$, $c_{i}=\|\mu_{i}^{(0)}\|_{2}^{2}$.
To guarantee convergence of the EM algorithm, we adopt the following assumption for the initialization parameters, consistent with the approach in [24].
  1. (A1)
    Suppose the GMM has parameters $\{(\pi_{j}^{\ast},\mu_{j}^{\ast}):j\in[K]\}$ such that
    $R_{\min}\geq C\cdot\sqrt{\log(\rho_{\pi}K)},$
    and suppose the mean initialization $\mu_{1}^{(0)},...,\mu_{K}^{(0)}$ satisfies
    $\forall i\in[K],\left\|{\mu_{i}^{(0)}-\mu_{i}^{\ast}}\right\|\leq\frac{R_{\min}}{16}.$
    Also, suppose the mixing weights are initialized such that
    $\forall i\in[K],\left|\pi_{i}^{(0)}-\pi_{i}^{\ast}\right|\leq\pi_{i}/2.$
We denote the output of the transformer ${\rm TF}_{\boldsymbol{\Theta}}$ as ${\boldsymbol{\theta}}^{\tt TF}\mathrel{\mathop{:}}={\{\pi_{1}^{\tt TF},{\mu}_{1}^{\tt TF},{\pi}_{2}^{\tt TF},\mu_{2}^{\tt TF},\cdots\pi_{K}^{\tt TF},\mu_{K}^{\tt TF}\}}$ and assume matched indices. Define
$D_{\boldsymbol{\Theta}}^{\tt TF}\mathrel{\mathop{:}}=\max_{i\in[K]}{\left\{\left\|{\mu_{i}^{\tt TF}-\mu_{i}^{\ast}}\right\|\vee\left(\left|\pi_{i}^{\tt TF}-\pi_{i}^{\ast}\right|/\pi_{i}\right)\right\}}.$
Now we propose the theorem that transformer can efficient approximate the EM Algorithm (Algorithm A.1), which is the formal version of Theorem 1.
Theorem C.1.
Fix $10<\delta,\beta<$ 1 and $11/2<a<$ 1. Suppose there exists a sufficiently large universal constant $C\geq 128$ for which assumption (A1) holds. If $N$ is suffcient large and $\varepsilon\leq 1/{\left(100K_{0}\right)}$ sufficient small such that
$\frac{\tilde{c}_{1}}{(1-a)\pi_{\min}}\sqrt{\frac{R_{\max}(R_{\max}\vee d)\log{\left(\frac{24K}{\delta}\right)}}{N}}+\tilde{c}_{2}\left(R_{\max}+d\left(1+\sqrt{\frac{2\log(\frac{4N}{\delta})}{d}}\right)\right)N\varepsilon<\frac{1}{2}\left(a-\frac{1}{2}\right),$
and
$\epsilon(N,\varepsilon,\delta,a)\mathrel{\mathop{:}}=\frac{\tilde{c}_{3}}{\left(1-a\right)\pi_{\min}}\sqrt{\frac{Kd\log(\frac{\tilde{C}N}{\delta})}{N}}+\tilde{c}_{4}{\left(\frac{1}{\pi_{\min}}+N{\left(R_{\max}+d+\sqrt{2d\log\left(\frac{4N}{\delta}\right)}\right)}\right)}\varepsilon<a(1-\beta),$
hold, where $\tilde{c}_{1}$-$\tilde{c}_{4}$ are universal constants, $\tilde{C}=288K^{2}(\sqrt{d}+2R_{\max}+\frac{1}{1-a})^{2}$. Then there exists a $2(L+1)$-layer transformer ${\rm TF}_{\boldsymbol{\Theta}}$ such that
$D_{\boldsymbol{\Theta}}^{\tt TF}\leq a\beta^{L}+\frac{1}{1-\beta}\epsilon(N,\varepsilon,\delta,a)$ (6)
holds with probability at least $1-\delta$. Moreover, ${\rm TF}_{\boldsymbol{\Theta}}$ falls within the class ${\mathcal{F}}$ with parameters satisfying:
$D=O(d_{0}+K_{0}),D^{\prime}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\varepsilon^{-1}\right)},M=O(1),\log B_{{\boldsymbol{\Theta}}}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\right)}.$
Notably, Eq. 6 holds for all tasks satisfying $d\leq d_{0}$ and $K\leq K_{0}$, where the parameters of transformer ${\boldsymbol{\Theta}}$ remains fixed across different tasks ${\mathcal{T}}$.
Remark C.1.
From Theorem C.1, if we take $\varepsilon=\tilde{O}{\left(N^{-3/2}d^{-1/2}\right)}$ and $L=O{\left(\log N\right)}$, then we have
$D_{\boldsymbol{\Theta}}^{\tt TF}\leq\tilde{O}{\left(\sqrt{\frac{d}{N}}\right)},$
which matches the canonical parametric error rate.
Remark C.2.
We give some explanations for the notations in Theorem C.1. Define
$D_{j}^{\rm pEM}\mathrel{\mathop{:}}=\max_{i\in[K]}{\left\{\left\|{\tilde{\mu}_{i}^{(j)}-\mu_{i}}\right\|\vee\left(\left|\tilde{\pi}_{i}^{(j)}-\pi_{i}\right|/\pi_{i}\right)\right\}},$
where ${\{\tilde{\mu}_{i}^{(j)},\tilde{\pi}_{i}^{(j)}\}}_{i\in[K]}$ are the parameters obtained at the $j$-th iteration of the population-EM algorithm (see Section C.3 for details). In the convergence analysis of the population-EM algorithm [24], it is shown that after the first iteration, the parameters lie in a small neighborhood of the true parameters with high probability (i.e., $D_{1}^{\rm pEM}\leq a$ for some $11/2\leq a<$ 1). Furthermore, the authors prove that the algorithm achieves linear convergence (i.e., $D_{j+1}^{\rm pEM}\leq\beta D_{j}^{\rm pEM}$ for $j\in\mathbb{N}_{+}$ and some $10<\beta<$ 1) with high probability if $D_{1}^{\rm pEM}\leq a$ holds. Following their notations, here $a$ represents the radius of the neighborhood after the first iteration, while $\beta$ is the linear convergence rate parameter. Finally, $\varepsilon$ controls the approximation error of the transformer.
C.2 Construction of transformer architecture and formal version of Fig. 7
In this section, we give the transformer architecture construction in Theorem C.1. We denote $w_{ij}=w_{j}(X_{i}),i\in[N],k\in[K]$ in this subsection for simplicity. Recall that we have assumed that $d\leq d_{0}$, $K\leq K_{0}$ and $N$ is divisible by $K$($N/K\in\mathbb{N}$). We first restate the encoding formulas in Eq. 5:
${\mathbf{H}}=\begin{bmatrix}\overline{X}_{1}&\overline{X}_{2}&\dots&\overline{X}_{N}\\ \overline{{\boldsymbol{\theta}}}_{1}&\overline{{\boldsymbol{\theta}}}_{2}&\dots&\overline{{\boldsymbol{\theta}}}_{N}\\ {\mathbf{p}}_{1}&{\mathbf{p}}_{2}&\dots&{\mathbf{p}}_{N}\end{bmatrix}\in\mathbb{R}^{D\times N},~{}\overline{{\boldsymbol{\theta}}}_{i}=\begin{bmatrix}\overline{{\boldsymbol{\pi}}}_{\log}\\ \overline{\mu}_{i\%K}\\ c_{i\%K}\\ \overline{{\mathbf{w}}}_{i}\\ \overline{{\mathbf{w}}}_{i\log}\\ \overline{{\boldsymbol{\pi}}}\end{bmatrix}\in\mathbb{R}^{d_{0}+4K_{0}+1},~{}{\mathbf{p}}_{i}\mathrel{\mathop{:}}=\begin{bmatrix}{\mathbf{0}}_{D-(2d_{0}+5K_{0}+2)}\\ 1\\ {\mathbf{e}}_{i\%K}\end{bmatrix}\in\mathbb{R}^{D-(2d_{0}+3K_{0}+1)},$ (7)
where $\overline{X}_{i}=[{X}_{i}^{\top},{\mathbf{0}}_{d_{0}-d}^{\top}]^{\top}$, $\overline{{\boldsymbol{\pi}}}_{\log}=[{\boldsymbol{\pi}}_{\log}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $\overline{\mu}_{i\%K}=[\mu_{i\%K}^{\top}$, ${\mathbf{0}}_{d_{0}-d}^{\top}]^{\top}$, $\overline{{\mathbf{w}}}_{i}=[{\mathbf{w}}_{i}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $\overline{{\mathbf{w}}}_{i\log}=[{\mathbf{w}}_{i\log}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $\overline{{\boldsymbol{\pi}}}=[{\boldsymbol{\pi}}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $c_{i\%K}\in\mathbb{R}$ and ${\mathbf{e}}_{i\%K}\in\mathbb{R}^{K_{0}}$ denotes the ${i\%K}$-th standard unit vector. For the initialization, we choose ${\boldsymbol{\pi}}_{\log}=\log{\boldsymbol{\pi}}^{(0)}$, $\mu_{i}=\mu_{i}^{(0)}$, $c_{i}=\|\mu_{i}^{(0)}\|_{2}^{2}$. and ${{\boldsymbol{\pi}}}={\mathbf{w}}_{i}={\mathbf{w}}_{i\log}={\mathbf{0}}_{K}$, $i\in[K]$. Finally, take ${\mathbf{H}}^{(0)}={\mathbf{H}}$ which is defined in Eq. 7.
Then in E-step, we consider the following attention structures: we define matrices ${\mathbf{Q}}^{(1)}$, ${\mathbf{K}}^{(1)}$, ${\mathbf{V}}^{(1)}$, such that
${\mathbf{Q}}^{(1)}{\mathbf{h}}_{i}^{(0)}=\begin{bmatrix}\overline{X}_{i}\\ \overline{{\boldsymbol{\pi}}}_{\log}\\ 1\\ {\mathbf{0}}\end{bmatrix},\quad{\mathbf{K}}^{(1)}{\mathbf{h}}_{j}^{(0)}=\begin{bmatrix}-\overline{\mu}_{j\%K}\\ {\mathbf{e}}_{j\%K}\\ \frac{1}{2}c_{j\%K}\\ {\mathbf{0}}\end{bmatrix},\quad{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{(0)}=\begin{bmatrix}{\mathbf{0}}_{d_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{d_{0}+1}\\ {\mathbf{e}}_{j\%K}\\ {\mathbf{0}}_{D-(2d_{0}+2K_{0}+1)}\end{bmatrix},$
and use the standard softmax attention, thus
$\widetilde{{\mathbf{h}}}_{i}^{(1)}={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(1)}}({\mathbf{H}}^{(0)})\right]}_{:,i}$ $={\mathbf{h}}_{i}^{(0)}+\sum_{j=1}^{N}{\left[\text{SoftMax}{\left({\left({\left({\mathbf{Q}}^{(1)}{\mathbf{h}}_{i}^{(0)}\right)}^{\top}{\left({\mathbf{K}}^{(1)}{\mathbf{h}}_{j}^{(0)}\right)}\right)}_{j=1}^{N}\right)}\right]}_{j}\cdot{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{(0)}$
$={\mathbf{h}}_{i}^{(0)}+\sum_{j=1}^{N}\frac{\alpha_{j\%K}^{(0)}\exp{\left(-X_{i}^{\top}\mu_{j\%K}+\frac{1}{2}\mu_{j\%K}^{\top}\mu_{j\%K}\right)}}{B\sum_{k=1}^{K}\alpha_{k}^{(0)}\exp{\left(-X_{i}^{\top}\mu_{k}+\frac{1}{2}\mu_{k}^{\top}\mu_{k}\right)}}\cdot{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{(0)}$
$={\mathbf{h}}_{i}^{(0)}+\frac{1}{B}\sum_{j=1}^{N}\hat{w}_{i\,j\%K}^{(1)}{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{(0)}$
$={\mathbf{h}}_{i}^{(0)}+\sum_{j=1}^{K}\hat{w}_{ij}^{(1)}{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{(0)}$
$={\mathbf{h}}_{i}^{(0)}+\begin{bmatrix}{\mathbf{0}}_{d_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{d_{0}+1}\\ \overline{\hat{{\mathbf{w}}}}_{i}^{(1)}\\ {\mathbf{0}}_{D-(2d_{0}+2K_{0}+1)}\end{bmatrix},~{}i\in[N].$
where $\overline{\hat{{\mathbf{w}}}}_{i}^{(1)}={\left(\hat{w}_{i1}^{(1)},\hat{w}_{i2}^{(1)},\cdots,\hat{w}_{iK}^{(1)},0,\cdots,0\right)}^{\top}\in\mathbb{R}^{K_{0}}$.
Then we use a two-layer MLP to approximate $\log x$ and clean all $\overline{{\boldsymbol{\pi}}}_{\log}$ , $\overline{\mu}_{i\%K}$ and $c_{i\%K}$, which is
${\mathbf{h}}_{i}^{(1)}=\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(1)}}{\left(\widetilde{{\mathbf{h}}}_{i}^{(1)}\right)}=\begin{bmatrix}\overline{X}_{i}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{d_{0}+1}\\ {\overline{\hat{{\mathbf{w}}}}}_{i}^{(1)}\\ {\overline{\hat{{\mathbf{w}}}}}_{i\log}^{(1)}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{D-(2d_{0}+5K_{0}+2)}\\ 1\\ {\mathbf{e}}_{i\%K}\end{bmatrix}~{},i\in[N],$
where ${\overline{\hat{{\mathbf{w}}}}}_{i\log}^{(1)}=\widehat{\log}{\overline{\hat{{\mathbf{w}}}}}_{i}^{(1)}$. Notice that although $\log x$ is not defined at 00, the MLP approximation is well defined with some value which we do not care because we will not use it in the M-step. Similarly, for any %$\ell\%2=1$, $\ell\in\mathbb{N}_{+}$ we have
${\mathbf{h}}_{i}^{(\ell)}=\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(\ell\%2)}}{\left({\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(\ell\%2)}}({\mathbf{H}}^{(\ell-1)})\right]}_{:,i}\right)}=\begin{bmatrix}\overline{X}_{i}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{d_{0}+1}\\ {\overline{\hat{{\mathbf{w}}}}}_{i}^{((\ell+1)/2)}\\ {\overline{\hat{{\mathbf{w}}}}}_{i\log}^{((\ell+1)/2)}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{D-(2d_{0}+5K_{0}+2)}\\ 1\\ {\mathbf{e}}_{i\%K}\end{bmatrix},~{}i\in[N],$
where ${\overline{\hat{{\mathbf{w}}}}}_{i\log}^{((\ell+1)/2)}=\widehat{\log}{\overline{\hat{{\mathbf{w}}}}}_{i}^{((\ell+1)/2)}$.
In M-step, we consider the following attention structures: we similarly define matrices ${\mathbf{Q}}_{m}^{(2)}$, ${\mathbf{K}}_{m}^{(2)}$, ${\mathbf{V}}_{m}^{(2)}$, $m=1,2$ such that
${\mathbf{Q}}_{1}^{(2)}{\mathbf{h}}_{j}^{(1)}=\begin{bmatrix}{\mathbf{e}}_{j\%K}\\ {\mathbf{0}}\end{bmatrix},\quad{\mathbf{K}}_{1}^{(2)}{\mathbf{h}}_{i}^{(1)}=\begin{bmatrix}{\overline{\hat{{\mathbf{w}}}}}_{i\log}^{(1)}\\ {\mathbf{0}}\end{bmatrix},\quad{\mathbf{V}}_{1}^{(2)}{\mathbf{h}}_{i}^{(1)}=\begin{bmatrix}{\mathbf{0}}_{d_{0}}\\ {\mathbf{0}}_{K_{0}}\\ \overline{X}_{i}\\ 0\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{D-(2d_{0}+3K_{0}+1)}\end{bmatrix},$
and
${\mathbf{Q}}_{2}^{(2)}{\mathbf{h}}_{j}^{(1)}={\mathbf{0}},\quad{\mathbf{K}}_{2}^{(2)}{\mathbf{h}}_{i}^{(1)}={\mathbf{0}},\quad{\mathbf{V}}_{2}^{(2)}{\mathbf{h}}_{i}^{(1)}=\begin{bmatrix}{\mathbf{0}}_{d_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{d_{0}+1}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\overline{\hat{{\mathbf{w}}}}}_{i}^{(1)}\\ {\mathbf{0}}_{D-(2d_{0}+4K_{0}+1)}\end{bmatrix},$
Then we get
$\widetilde{{\mathbf{h}}}_{j}^{(2)}={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(2)}}({\mathbf{H}}^{(1)})\right]}_{:,j}$ $={\mathbf{h}}_{j}^{(1)}+\sum_{m=1}^{2}\sum_{i=1}^{N}{\left[\text{SoftMax}{\left({\left({\left({\mathbf{Q}}_{m}^{(2)}{\mathbf{h}}_{i}^{(1)}\right)}^{\top}{\left({\mathbf{K}}_{m}^{(2)}{\mathbf{h}}_{j}^{(1)}\right)}\right)}_{j=1}^{N}\right)}\right]}_{j}\cdot{\mathbf{V}}_{m}^{(2)}{\mathbf{h}}_{i}^{(1)}$
$={\mathbf{h}}_{j}^{(1)}+\sum_{i=1}^{N}\frac{\hat{w}_{i\,j\%K}^{(1)}}{\sum_{i=1}^{N}\hat{w}_{i\,j\%K}^{(1)}}\cdot{\mathbf{V}}_{1}^{(2)}{\mathbf{h}}_{i}^{(1)}+\sum_{i=1}^{N}\frac{1}{N}\cdot{\mathbf{V}}_{2}^{(2)}{\mathbf{h}}_{i}^{(1)}$
$={\mathbf{h}}_{j}^{(1)}+\begin{bmatrix}{\mathbf{0}}_{d_{0}}\\ {\mathbf{0}}_{K_{0}}\\ \overline{\hat{\mu}}_{j\%K}^{(1)}\\ 0\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{D-(2d_{0}+3K_{0}+1)}\end{bmatrix}+\begin{bmatrix}{\mathbf{0}}_{d_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{d_{0}+1}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{K_{0}}\\ \overline{\hat{{\boldsymbol{\pi}}}}^{(1)}\\ {\mathbf{0}}_{D-(2d_{0}+4K_{0}+1)}\end{bmatrix},$
$={\mathbf{h}}_{j}^{(1)}+\begin{bmatrix}{\mathbf{0}}_{d_{0}}\\ {\mathbf{0}}_{K_{0}}\\ \overline{\hat{\mu}}_{j\%K}^{(1)}\\ 0\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{K_{0}}\\ \overline{\hat{{\boldsymbol{\pi}}}}^{(1)}\\ {\mathbf{0}}_{D-(2d_{0}+4K_{0}+1)}\end{bmatrix},~{}j\in[N].$
Similarly, we use a two-layer MLP to approximate $\log x$, $x^{2}$ and clean all $\overline{{\mathbf{w}}}_{i}$, $\overline{{\mathbf{w}}}_{i\log}$ and $\overline{{\boldsymbol{\pi}}}_{i}$, which is
${\mathbf{h}}_{j}^{(2)}=\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(2)}}{\left(\widetilde{{\mathbf{h}}}_{i}^{(2)}\right)}=\begin{bmatrix}\overline{X}_{j}\\ \overline{\hat{{\boldsymbol{\pi}}}}_{\log}^{(1)}\\ \overline{\hat{\mu}}_{j\%K}^{(1)}\\ \hat{c}_{j\%K}^{(1)}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{D-(2d_{0}+5K_{0}+2)}\\ 1\\ {\mathbf{e}}_{j\%K}\end{bmatrix}~{},j\in[N],$
where $\overline{\hat{{\boldsymbol{\pi}}}}_{\log}^{(1)}=\widehat{\log}\overline{\hat{{\boldsymbol{\pi}}}}^{(1)}$, $\hat{c}_{j\%K}^{(1)}=\widehat{\|\hat{\mu}_{j\%K}^{(1)}\|_{2}^{2}}$.
Similarly, for any %$\ell\%2=0$, $\ell\in\mathbb{N}_{+}$ we have
${\mathbf{h}}_{j}^{(\ell)}=\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(\ell\%2)}}{\left({\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(\ell\%2)}}({\mathbf{H}}^{(\ell-1)})\right]}_{:,j}\right)}=\begin{bmatrix}\overline{X}_{j}\\ \overline{\hat{{\boldsymbol{\pi}}}}_{\log}^{(\ell/2)}\\ \overline{\hat{\mu}}_{j\%K}^{(\ell/2)}\\ \hat{c}_{j\%K}^{(\ell/2)}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{D-(2d_{0}+5K_{0}+2)}\\ 1\\ {\mathbf{e}}_{j\%K}\end{bmatrix}~{},j\in[N].$
where $\overline{\hat{{\boldsymbol{\pi}}}}_{\log}^{(\ell/2)}=\widehat{\log}\overline{\hat{{\boldsymbol{\pi}}}}^{(\ell/2)}$, $\hat{c}_{j\%K}^{(\ell/2)}=\widehat{\|\hat{\mu}_{j\%K}^{(\ell/2)}\|_{2}^{2}}$.
Thus, we can get $\hat{{\boldsymbol{\pi}}}^{(\ell)}$ and $\hat{\mu}_{j}^{(\ell)}$, $j\in[K]$ after $2\ell$ layers of transformer constructed above. (The last-layer MLP block retains ${\boldsymbol{\pi}}$ as an output parameter without cleaning it.) Our transformer construction is summarized in Fig. 9, which is the formal version of Fig. 7 in Section 4.2.
Remark C.3.
The output of transformer ${\mathbf{H}}^{(2L)}$ is a large matrix containing lots of elements. To get the estimated parameters, we need to extract specific elements. In details, ${\mathbf{H}}^{(2L)}={\left[{\mathbf{h}}_{1}^{(2L)},\cdots,{\mathbf{h}}_{N}^{(2L)}\right]}\in\mathbb{R}^{D\times N}$, where
${\mathbf{h}}_{i}^{(2L)}=\begin{bmatrix}\overline{X}_{i}\\ \overline{\hat{{\boldsymbol{\pi}}}}_{\log}^{(L)}\\ \overline{\hat{\mu}}_{i\%K}^{(L)}\\ \hat{c}_{i\%K}^{(L)}\\ {\mathbf{0}}_{K_{0}}\\ {\mathbf{0}}_{K_{0}}\\ \overline{\hat{{\boldsymbol{\pi}}}}^{(L)}\\ {\mathbf{0}}_{D-(2d_{0}+5K_{0}+2)}\\ 1\\ {\mathbf{e}}_{j\%K}\end{bmatrix}~{},i\in[N].$
We use the following linear attentive pooling to get the parameters:
$\mathbf{O}=\frac{1}{N}({\mathbf{V}}_{o}{\mathbf{H}}){\left(({\mathbf{K}}_{o}{\mathbf{H}})^{\top}{\mathbf{Q}}_{o}\right)}\in\mathbb{R}^{(d+K)\times K},$
where ${\mathbf{Q}}_{o}={\left[{\bf q}_{o1},\cdots,{\bf q}_{oK}\right]}\in\mathbb{R}^{(d+K)\times K}$, ${\mathbf{K}}_{o},{\mathbf{V}}_{o}\in\mathbb{R}^{(d+K)\times N}$ satisfying
${\bf q}_{oi}=\begin{bmatrix}K{\mathbf{e}}_{i}\\ {\mathbf{0}}_{d}\end{bmatrix},~{}{\mathbf{K}}_{o}{\mathbf{h}}_{j}^{(2L)}=\begin{bmatrix}{\mathbf{e}}_{j\%K}\\ {\mathbf{0}}_{d}\end{bmatrix},~{}{\mathbf{V}}_{o}{\mathbf{h}}_{j}^{(2L)}=\begin{bmatrix}\hat{{\boldsymbol{\pi}}}^{(L)}\\ \hat{\mu}_{j\%K}\end{bmatrix}.$
Thus by $N/K\in\mathbb{N}$, we have
$\mathbf{o}_{i}=\frac{1}{N}\sum_{j\in[N]}{\bf q}_{oi}^{\top}{\left({\mathbf{K}}_{o}{\mathbf{h}}_{j}\right)}{\mathbf{V}}_{o}{\mathbf{h}}_{j}=\frac{K}{N}\frac{N}{K}\begin{bmatrix}\hat{{\boldsymbol{\pi}}}^{(L)}\\ \hat{\mu}_{i}\end{bmatrix}=\begin{bmatrix}\hat{{\boldsymbol{\pi}}}^{(L)}\\ \hat{\mu}_{i}\end{bmatrix}\in\mathbb{R}^{(d+K)},i\in[K].$
Finally, we get
$\mathbf{O}={\left[{\bf q}_{o1},\cdots,{\bf q}_{oN}\right]}=\begin{bmatrix}\hat{{\boldsymbol{\pi}}}^{(L)}&\hat{{\boldsymbol{\pi}}}^{(L)}&\cdots&\hat{{\boldsymbol{\pi}}}^{(L)}\\ \hat{\mu}_{1}&\hat{\mu}_{2}&\cdots&\hat{\mu}_{K}\end{bmatrix}.$
Figure 9: Transformer Construction for Approximating EM Algorithm Iterations. The pink box represents the state of tokens, while the blue box represents the structure of different parts of the network. The term "clean" means setting all positions of the corresponding vector to zero.
C.3 Convergence results for EM algorithm
C.3.1 Convergence results for population-EM algorithm
First, we review some notations. Recall that $\pi_{min}=\min_{i}\pi_{i}^{*}$, $\rho_{\pi}=\max_{i}\pi_{i}^{*}/\min_{i}\pi_{i}^{*}$, $R_{ij}=\|\mu_{i}^{*}-\mu_{j}^{*}\|$, $R_{min}=\min_{i\neq j}R_{ij}$ and $R_{max}={\left(\max_{i\neq j}R_{ij}\right)}\vee{\left(\max_{i\in[K]}\left\|{\mu_{i}^{\ast}}\right\|\right)}$. Without the loss of generality, we assume that $R_{max}\geq 1$. For clarity, we restate assumption ((A1)), which is consistent with [24].
  1. (A1)
    Suppose the GMM has parameters $\{(\pi_{j}^{\ast},\mu_{j}^{\ast}):j\in[K]\}$ such that
    $R_{\min}\geq C\cdot\sqrt{\log(\rho_{\pi}K)},$ (8)
    and suppose the mean initialization $\mu_{1}^{(0)},...,\mu_{K}^{(0)}$ satisfies
    $\forall i\in[K],\left\|{\mu_{i}^{(0)}-\mu_{i}^{\ast}}\right\|\leq\frac{R_{\min}}{16}.$ (9)
    Also, suppose the mixing weights are initialized such that
    $\forall i\in[K],\left|\pi_{i}^{(0)}-\pi_{i}^{\ast}\right|\leq\pi_{i}/2.$ (10)
For population-EM, the algorithm can be presented as
(E-step): $w_{i}(X)$ $=\frac{\pi_{i}\exp(-\|X-\mu_{i}\|^{2}/2)}{\sum_{j=1}^{K}\pi_{j}\exp(-\|X-\mu_{j}\|^{2}/2)},$
(M-step): $\pi_{i}^{+}$ $=\mathbb{E}[w_{i}],\quad\mu_{i}^{+}=\mathbb{E}[w_{i}X]/\mathbb{E}[w_{i}].$
The following results gives linear convergenve guarantees of population-EM, which comes from [24].
Theorem C.2 ([24], Theorem 1, part i).
Let $C\geq 64$ be a universal constant for which assumption ((A1)) holds. Then, after one-step population-EM update, we have
$\forall i\in[K],|\pi_{i}^{+}-\pi_{i}^{*}|\leq\pi_{i}^{*}/2,~{}\|\mu_{i}^{+}-\mu_{i}^{*}\|\leq 1/2.$ (11)
Now we define
$D_{m}=\max_{i\in[K]}\left(\|\mu_{i}-\mu_{i}^{*}\|\vee|\pi_{i}-\pi_{i}^{*}|/\pi_{i}^{*}\right),$
and
$D_{m}^{+}=\max_{i\in[K]}\left(\|\mu_{i}^{+}-\mu_{i}^{*}\|\vee|\pi_{i}^{+}-\pi_{i}^{*}|/\pi_{i}^{*}\right).$
The linear convergence of population-EM is stated by the following theorem.
Theorem C.3 ([24], Theorem 1, part ii).
Let $C\geq 128$ be a large enough universal constant. Fix $10<a<$ 1. Suppose the separation condition Eq. 8 holds and suppose the initialization parameter satisfies $D_{m}\leq a$, then $D_{m}^{+}\leq\beta D_{m}$ for some $10<\beta<$ 1.
Remark C.4.
Here the contraction parameter $\beta$ is only dependent with $C$ and $a$. In other words, if we fix $a\in(0,1)$, then for any $\beta\in(0,1)$, there exists a large enough $C$ such that Theorem C.3 holds. For details, see Appendix E in [24].
Combing Theorem C.2 and Theorem C.3, we can get the linear convergence of population-EM algorithm.
C.3.2 Convergence results for empirical-EM algorithm
Now we consider the empirical-EM, i.e., Algorithm A.1. For convenience, the algorithm can be presented as
(E-step): $w_{i}(X_{\ell})$ $=w_{\ell i}=\frac{\pi_{i}\exp(-\|X_{\ell}-\mu_{i}\|^{2}/2)}{\sum_{j=1}^{K}\pi_{j}\exp(-\|X_{\ell}-\mu_{j}\|^{2}/2)}$
(M-step): $\pi_{i}^{+}$ $=\frac{1}{n}\sum_{l=1}^{n}w_{i}(X_{\ell}),~{}\mu_{i}^{+}=\frac{\sum_{l=1}^{n}w_{i}(X_{\ell})X_{\ell}}{\sum_{l=1}^{n}w_{i}(X_{\ell})}=\frac{1}{n\pi_{i}^{+}}\sum_{l=1}^{n}w_{i}(X_{\ell})X_{\ell}.$
Similarly, we can define $D_{m}$ and $D_{m}^{+}$ in empirical sense.
For the linear convergence of empirical-EM, we have the following theorem.
Theorem C.4.
Fix $10<\delta,\beta<$ 1 and $10<a<$ 1. Let $C\geq 128$ be a large enough universal constant. Suppose the separation condition Eq. 8 holds and suppose the initialization parameter satisfies $D_{m}\leq a$. If $n$ is suffcient large such that
$\varepsilon_{unif}:=\frac{\tilde{c}}{\left(1-a\right)\pi_{min}}\sqrt{\frac{Kd\log(\frac{\tilde{C}n}{\delta})}{n}}<a(1-\beta)$
where $\tilde{C}=72K^{2}(\sqrt{d}+2R_{max}+\frac{1}{1-a})^{2}$ and $\tilde{c}$ is a universal constant. Then
$D_{m}^{+}\leq\beta D_{m}+\varepsilon_{unif}\leq a$
uniformly holds with probability at least $1-\delta$.
Proof.
First, we have
$\frac{|\pi_{i}^{+}-\pi_{i}^{*}|}{\pi_{i}^{*}}$ $=\frac{1}{\pi_{i}^{*}}\left|\frac{1}{n}\sum_{l=1}^{n}w_{i}(X_{\ell})-\pi_{i}^{*}\right|$
$\leq\frac{1}{\pi_{i}^{*}}\left(\left|\frac{1}{n}\sum_{l=1}^{n}w_{i}(X_{\ell})-\mathbb{E}\left[w_{i}(X)\right]\right|+\left|\mathbb{E}\left[w_{i}(X)\right]-\pi_{i}^{*}\right|\right)$
$:=(I)+(II).$
By Theorem C.3, we get
$(II)=\frac{1}{\pi_{i}^{*}}\left|\mathbb{E}\left[w_{i}(X)\right]-\pi_{i}^{*}\right|\leq\tilde{\beta}D_{m}.$
And by Lemma C.2, we have
$(I)=\frac{1}{\pi_{i}^{*}}\left|\frac{1}{n}\sum_{l=1}^{n}w_{i}(X_{\ell})-\mathbb{E}\left[w_{i}(X)\right]\right|\leq\frac{\tilde{c}_{1}}{\pi_{min}}\sqrt{\frac{Kd\log(\frac{\tilde{C}_{1}n}{\delta_{1}})}{n}},$
where $\tilde{C}_{1}=18K^{2}(\sqrt{d}+2R_{max}+\frac{1}{1-a})$ and $\tilde{c}_{1}$ is a suitable universal constant. Thus, by taking $\tilde{\beta}=\beta$. $\delta_{1}=\delta/2$ and suitable $\tilde{c}$, ${|\pi_{i}^{+}-\pi_{i}^{*}|}/{\pi_{i}^{*}}\leq\beta D_{m}+\varepsilon_{unif}\leq a$, $\forall i\in[K]$.
For the second term, we have
$\|\mu_{i}^{+}-\mu_{i}^{*}\|$ $=\left\|\frac{1}{n\pi_{i}^{+}}\sum_{l=1}^{n}w_{i}(X_{\ell})(X_{\ell}-\mu_{i}^{\ast})\right\|$
$\leq\frac{1}{\pi_{i}^{+}}{\left(\left\|\frac{1}{n}\sum_{l=1}^{n}w_{i}(X_{\ell})(X_{\ell}-\mu_{i}^{\ast})-\mathbb{E}{\left[w_{i}(X)(X-\mu_{i}^{\ast})\right]}\right\|+\left\|\mathbb{E}{\left[w_{i}(X)(X-\mu_{i}^{\ast})\right]}\right\|\right)}$
$:=(III)+(IV),$
By Theorem C.3 and Remark C.4 we get,
$(IV)$ $=\frac{1}{\pi_{i}^{+}}\left\|\mathbb{E}{\left[w_{i}(X)(X-\mu_{i}^{\ast})\right]}\right\|$
$\overset{(i)}{\leq}\frac{1}{(1-a)\pi_{i}^{\ast}}\left\|\mathbb{E}{\left[w_{i}(X)(X-\mu_{i}^{\ast})\right]}\right\|$
$=\frac{\mathbb{E}{\left[w_{i}(X)\right]}}{(1-a)\pi_{i}^{\ast}}\left\|\frac{\mathbb{E}{\left[w_{i}(X)X\right]}}{\mathbb{E}{\left[w_{i}(X)\right]}}-\mu_{i}^{\ast}\right\|$
$\leq\frac{1+a}{1-a}\tilde{\beta}D_{m}.$
where $(i)$ follows from ${|\pi_{i}^{+}-\pi_{i}^{*}|}/{\pi_{i}^{*}}\leq a$. And by Lemma C.3, we have
$(III)$ $=\frac{1}{\pi_{i}^{+}}{\left(\left\|\frac{1}{n}\sum_{l=1}^{n}w_{i}(X_{\ell})(X_{\ell}-\mu_{i}^{\ast})-\mathbb{E}{\left[w_{i}(X)(X-\mu_{i}^{\ast})\right]}\right\|\right)}$
$\leq\frac{1}{(1-a)\pi_{i}^{\ast}}{\left(\left\|\frac{1}{n}\sum_{l=1}^{n}w_{i}(X_{\ell})(X_{\ell}-\mu_{i}^{\ast})-\mathbb{E}{\left[w_{i}(X)(X-\mu_{i}^{\ast})\right]}\right\|\right)}$
$\leq\frac{\tilde{c}_{2}}{(1-a)\pi_{min}}\sqrt{\frac{Kd\log(\frac{\tilde{C}_{2}n}{\delta_{2}})}{n}},$
where $\tilde{C}=18K^{2}(\sqrt{d}+2R_{max}+\frac{1}{1-a})^{2}$ and $\tilde{c}$ is a suitable universal constant. Thus, by taking $\tilde{\beta}=(1-a)/(1+a)\beta$, $\delta_{2}=\delta/2$ and suitable $\tilde{c}$, $\|\mu_{i}^{+}-\mu_{i}^{*}\|\leq\beta D_{m}+\varepsilon_{unif}\leq a$, $\forall i\in[K]$.
In conclusion, if we take $\tilde{\beta}=(1-a)/(1+a)\beta$, $\tilde{c}=\tilde{c}_{1}\vee\tilde{c}_{2}$, $C=C(\beta,a)\geq 128$ large enough such that Theorem C.3 holds, and take $\delta_{1}=\delta_{2}=\delta/2$ and use union bound argument, then we get $D_{m}^{+}\leq\beta D_{m}+\varepsilon_{unif}\leq a$.
We need the following technical lemma.
Lemma C.1 ([40], Lemma B.2.).
Fix $10<\delta<$ 1. Let $B_{1},\ldots,B_{K}\subset\mathbb{R}^{d}$ be Euclidean balls of radii $r_{1},\ldots,r_{K}$. Define $\mathcal{B}=\otimes_{k=1}^{K}B_{k}\subset\mathbb{R}^{Kd}$ and $r=\max_{k\in[K]}r_{k}$. Let $X$ be a random vector in $\mathbb{R}^{d}$ and $W:\mathbb{R}^{d}\times{\cal B}\to\mathbb{R}^{k}$ where $k\leq d$. Assume the following hold:
1. There exists a constant $L\geq 1$ such that for any $\theta\in\mathcal{B},\varepsilon>0$, and $\theta^{\varepsilon}\in\mathcal{B}$ which satisfies $\max_{i\in[K]}\|\theta_{i}-\theta_{i}^{\varepsilon}\|\leq\varepsilon$, then $\mathbb{E}_{X}\left[\sup_{\mu\in\mathcal{B}}\|W(X,\theta)-W(X,\theta^{\varepsilon})\|\right]\leq L\varepsilon$.
2. There exists a constant $R$ such that for any $\theta\in{\cal{B}}$, $\|W(X,\theta)\|_{\psi_{2}}\leq R$.
Let $X_{1},\ldots,X_{n}$ be i.i.d. random vectors with the same distribution as $X$. Then there exists a universal constant $\tilde{c}$ such that with probability at least $1-\delta$,
$\sup_{\theta\in\mathcal{B}}\left\|\frac{1}{n}\sum_{\ell=1}^{n}W\left(X_{\ell},\theta\right)-\mathbb{E}\left[W\left(X,\theta\right)\right]\right\|\leq R\sqrt{\tilde{c}\frac{Kd\log\left(1+\frac{12nLr}{\delta}\right)}{n}}.$ (12)
Remark C.5.
There is one difference between Lemma C.1 and LemmaB.2. in [40]: in Lemma C.1, we use $1+\frac{12nLr}{\delta}$ to replace $\frac{18nLr}{\delta}$, thus we avoid the condition $r_{1},\cdots,r_{K}\geq 1$.
Hence we can get the uniform convergence of $w_{i}(X,\theta)$ and $w_{i}(X,\theta)(X-\mu_{i}^{*})$, $i\in[K]$. Our proof is similar to [40], except that we consider the variation of both $\pi$ and $\mu$. From now on, we denote $\theta_{i}=\{\pi_{i},\mu_{i}\}$, $\theta=\{\theta_{i}\}_{i=1}^{n}$.
Lemma C.2.
[…]
$\sup_{\theta\in{\mathcal{D}}_{a}}\left|\frac{1}{n}\sum_{\ell=1}^{n}w_{i}(X_{\ell},\theta)-\mathbb{E}[w_{i}(X,\theta)]\right|\leq\tilde{c}\sqrt{\frac{Kd\log(\frac{\tilde{C}n}{\delta})}{n}},~{}\forall i\in[K],$ (13)
where $\tilde{C}=18K^{2}(\sqrt{d}+2R_{max}+\frac{1}{1-a})$ and $\tilde{c}$ is a suitable universal constant.
Proof.
The proof is similar to the proof of Lemma 5.1 in [40]. For simplicity, we omit it. ∎
Lemma C.3.
[…]
$\sup_{\theta\in{\mathcal{D}}_{a}}\left|\frac{1}{n}\sum_{\ell=1}^{n}w_{i}(X_{\ell},\theta)(X_{\ell}-\mu_{i}^{*})-\mathbb{E}[w_{i}(X,\theta)(X_{\ell}-\mu_{i}^{*})]\right|\leq\tilde{c}\sqrt{\frac{Kd\log(\frac{\tilde{C}n}{\delta})}{n}},~{}\forall i\in[K],$ (14)
where $\tilde{C}=36K^{2}(\sqrt{d}+2R_{max}+\frac{1}{1-a})^{2}$ and $\tilde{c}$ is a suitable universal constant.
Proof.
The proof is similar to the proof of Lemma 5.4 in [40](Notice that the condition (36) in [40] is trivial in our case). For simplicity, we omit it. ∎
For the first step empirical-EM, we have the following results.
Theorem C.5.
Fix $10<\delta<$ 1 and $11/2<a<$ 1. Let $C\geq 128$ be a large enough universal constant for which assumption ((A1)) holds. If $n$ is suffcient large such that
$\varepsilon_{step1}:=\frac{\tilde{c}}{(1-a)\pi_{min}}\sqrt{\frac{R_{max}(R_{max}\vee d)\log{\left(\frac{6K}{\delta}\right)}}{n}}<\left(a-\frac{1}{2}\right),$
where $\tilde{c}$ is a universal constant. Then
$D_{m}^{+}\leq\frac{1}{2}+\varepsilon_{step1}\leq a$
holds with probability at least $1-\delta$.
Proof.
Notice that we only need simple concentration not uniform concentration in this theorem. We use the same definition of term $(I)$, $(II)$ as in the proof of Theorem C.4. First, by Theorem C.2, we have $(II)\leq 1/2$. Since $0\leq w_{i}(X)\leq 1$, by a standard concentration of bounded variables, we can get
$(I)\leq\frac{\tilde{c}_{1}}{\pi_{min}}\sqrt{\frac{\log(\frac{K}{\delta_{1}})}{n}},\forall i\in[K],$ (15)
where $\tilde{c}_{1}$ is a universal constant. Taking $\tilde{c}\geq\tilde{c}_{1}$ and $\delta_{1}=\delta/2$, we have
$\frac{|\pi_{i}^{+}-\pi_{i}^{*}|}{\pi_{i}^{*}}\leq\frac{1}{2}+\frac{\tilde{c}}{\pi_{min}}\sqrt{\frac{\log{\left(\frac{2K}{\delta}\right)}}{n}}\leq a,\forall i\in[K].$
For the second term, we have
$\|\mu_{i}^{+}-\mu_{i}^{*}\|$ $=\left\|\frac{1}{n\pi_{i}^{+}}\sum_{l=1}^{n}w_{i}(X_{\ell})X_{\ell}-\mu_{i}^{\ast}\right\|$
$\leq\left\|\frac{1}{n\pi_{i}^{+}}\sum_{l=1}^{n}w_{i}(X_{\ell})X_{\ell}-\frac{\mathbb{E}[w_{i}(X_{\ell})X_{\ell}]}{\mathbb{E}[w_{i}(X_{\ell})]}\right\|+\left\|\frac{\mathbb{E}[w_{i}(X_{\ell})X_{\ell}]}{\mathbb{E}[w_{i}(X_{\ell})]}-\mu_{i}^{\ast}\right\|$
$:=(V)+(VI).$
By Theorem C.2, we have $(VI)\leq 1/2$. For $(V)$, by triangle inequality,
$(V)$ $\leq\left\|\frac{1}{n\pi_{i}^{+}}\sum_{l=1}^{n}w_{i}(X_{\ell})X_{\ell}-\frac{1}{\pi_{i}^{+}}\mathbb{E}[w_{i}(X_{\ell})X_{\ell}]\right\|+\left\|\frac{1}{\pi_{i}^{+}}\mathbb{E}[w_{i}(X_{\ell})X_{\ell}]-\frac{\mathbb{E}[w_{i}(X_{\ell})X_{\ell}]}{\mathbb{E}[w_{i}(X_{\ell})]}\right\|$
$=\frac{1}{\pi_{i}^{+}}\left\|\frac{1}{n}\sum_{l=1}^{n}w_{i}(X_{\ell})X_{\ell}-\mathbb{E}[w_{i}(X_{\ell})X_{\ell}]\right\|+\frac{\left\|\mathbb{E}[w_{i}(X_{\ell})X_{\ell}]\right\|}{\pi_{i}^{+}\mathbb{E}[w_{i}(X_{\ell})]}\left|{\pi_{i}^{+}}-{\mathbb{E}[w_{i}(X_{\ell})]}\right|.$ (16)
Using Lemma B.1 and Lemma B.2 in [50], we can get $\|w_{i}(X_{\ell})X_{\ell}\|_{\psi_{2}}\leq\|X_{\ell}\|_{\psi_{2}}\leq\tilde{c}_{3}R_{max}$, $\forall i\in[K]$. Hence by Lemma B.1 in [40], with probability at least $1-\delta_{2}$,
$\left\|\frac{1}{n}\sum_{l=1}^{n}w_{i}(X_{\ell})X_{\ell}-\mathbb{E}[w_{i}(X_{\ell})X_{\ell}]\right\|\leq\tilde{c}_{4}\sqrt{\frac{R_{max}d\log{\left(\frac{3K}{\delta_{2}}\right)}}{n}},~{}\forall i\in[K],$
where $\tilde{c}_{4}$ is an universal constant. And by Theorem C.2, we have
$\frac{\left\|\mathbb{E}[w_{i}(X_{\ell})X_{\ell}]\right\|}{\mathbb{E}[w_{i}(X_{\ell})]}\leq R_{max}+\frac{1}{2}\leq 2R_{max},\forall i\in[K].$
Finally, by Eq. 15,
$\left|{\pi_{i}^{+}}-{\mathbb{E}[w_{i}(X_{\ell})]}\right|\leq{\tilde{c}_{1}}\sqrt{\frac{\log(\frac{K}{\delta_{1}})}{n}}.$
Combining all terms together and taking $\delta_{1}=\delta_{2}=\delta/2$ we can bound Eq. 16 by
$(V)$ $\leq\frac{1}{\pi_{i}^{+}}{\left(\tilde{c}_{4}\sqrt{\frac{R_{max}d\log{\left(\frac{3K}{\delta_{2}}\right)}}{n}}+2\tilde{c}_{1}R_{max}\sqrt{\frac{\log(\frac{K}{\delta_{1}})}{n}}\right)}$
$\leq\frac{\tilde{c}_{6}}{(1-a)\pi_{min}}\sqrt{\frac{R_{max}(R_{max}\vee d)\log{\left(\frac{6K}{\delta}\right)}}{n}}.$
Taking $\tilde{c}\geq\tilde{c}_{6}$, we get
$\|\mu_{i}^{+}-\mu_{i}^{*}\|\leq\frac{1}{2}+\frac{\tilde{c}}{(1-a)\pi_{min}}\sqrt{\frac{R_{max}(R_{max}\vee d)\log{\left(\frac{6K}{\delta}\right)}}{n}}\leq a.$
C.3.3 Convergence results for transformer-based EM in Section C.2
We first state some useful approximation lemmas.
Lemma C.4 (Lemma 9 in [33]).
For any $A>0$, $\delta>0$, take $M=\lceil 2{\log A}/\delta\rceil+1\in\mathbb{N}$. Then there exists $\{(a_{j},w_{j},b_{j})\}_{j\in[M]}$ with
$\sup_{j}|a_{j}|\leq 2A,~{}~{}~{}\sup_{j}|w_{j}|\leq 1,~{}~{}~{}\sup_{j}|b_{j}|\leq A,$ (17)
such that defining $\log_{\delta}:\mathbb{R}\to\mathbb{R}$ by
$\log_{\delta}(x)=\sum_{j=1}^{M}a_{j}\cdot\operatorname{ReLU}(w_{j}x+b_{j}),$
we have $\log_{\delta}$ is non-decreasing on $[1/A,A]$, and
$\sup_{x\in[1/A,A]}|\log(x)-\log_{\delta}(x)|\leq\delta.$
Remark C.6.
There is a small improvement $M=\lceil 2\log A/\delta\rceil+1$ compared to $M=\lceil 2A/\delta\rceil+1$ in [33]. Further more, it is easy to check that $\log_{\delta}(x)\leq-\log A$ for $x\in[0,1/A]$.
Lemma C.5.
For any $A>0$, $\delta>0$, take $M=\lceil 2A^{2}/\delta\rceil+1\in\mathbb{N}$. Then there exists $\{(a_{j},w_{j},b_{j})\}_{j\in[M]}$ with
$\sup_{j}|a_{j}|\leq 2A,~{}~{}~{}\sup_{j}|w_{j}|\leq 1,~{}~{}~{}\sup_{j}|b_{j}|\leq A,$ (18)
such that defining $\phi_{\delta}:\mathbb{R}\to\mathbb{R}$ by
$\phi_{\delta}(x)=\sum_{j=1}^{M}a_{j}\cdot\operatorname{ReLU}(w_{j}x+b_{j}),$
we have $\phi_{\delta}$ is non-decreasing on $[-A,A]$, and
$\sup_{x\in[-A,A]}|\phi_{\delta}(x)-x^{2}|\leq\delta.$
Proof.
Similar to Lemma C.4. Omitted. ∎
Lemma C.6 (Lemma A.1 in [4]).
Let $\beta\sim\mathcal{N}(0,I_{d})$. Then we have
$\mathbb{P}\Big{(}\|\beta\|^{2}\geq d(1+\delta)^{2}\Big{)}\leq e^{-d\delta^{2}/2}.$
Lemma C.7 (Lemma 18 in [28]).
For any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{d}$, we have
$\left\|\log\left(\frac{e^{\mathbf{u}}}{\|e^{\mathbf{u}}\|_{1}}\right)-\log\left(\frac{e^{\mathbf{v}}}{\|e^{\mathbf{v}}\|_{1}}\right)\right\|_{\infty}\leq 2\left\|\mathbf{u}-\mathbf{v}\right\|_{\infty}.$
Corollary C.1.
For any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{d}$, we have
$\left\|\frac{e^{\mathbf{u}}}{\|e^{\mathbf{u}}\|_{1}}-\frac{e^{\mathbf{v}}}{\|e^{\mathbf{v}}\|_{1}}\right\|_{\infty}\leq\exp{\left(2\left\|\mathbf{u}-\mathbf{v}\right\|_{\infty}\right)}-1$
Proof.
This follows directly from Lemma C.7 and simple calculations. ∎
Now we propose the results for transformer-based EM. Similar to Section C.2, we use notations with superscript “^” to represent the output of the transformer-based EM.
Theorem C.6.
Fix $10<\delta<$ 1 and $11/2<a<$ 1. Let $C\geq 128$ be a large enough universal constant for which assumption ((A1)) holds. If $n$ is sufficient large and $\varepsilon\leq 1/100$ sufficient small such that
$\frac{\tilde{c}_{1}}{(1-a)\pi_{min}}\sqrt{\frac{R_{max}(R_{max}\vee d)\log{\left(\frac{12K}{\delta}\right)}}{n}}+\tilde{c}_{2}\left(R_{\max}+d\left(1+\sqrt{\frac{2\log(\frac{2n}{\delta})}{d}}\right)\right)n\varepsilon<\frac{1}{2}\left(a-\frac{1}{2}\right),$
where $\tilde{c}_{1}$, $\tilde{c}_{2}$ are universal constants. Then there exists a 2-layer transformer ${\rm TF}_{\boldsymbol{\Theta}}$ such that $\hat{D}_{m}^{+}\leq a$ holds with probability at least $1-\delta$. Moreover, ${\rm TF}_{\boldsymbol{\Theta}}$ falls within the class ${\mathcal{F}}$ with parameters satisfying:
$D=O(d_{0}+K_{0}),D^{\prime}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\varepsilon^{-1}\right)},M=O(1),\log B_{{\boldsymbol{\Theta}}}\leq O{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\right)}.$
Proof.
Recall that $\hat{D}_{m}^{+}=\max_{i\in[K]}\left(\|\hat{\mu}_{i}^{+}-\mu_{i}^{*}\|\vee|\hat{\pi}_{i}^{+}-\pi_{i}^{*}|/\pi_{i}^{*}\right)$. Thus
$\hat{D}_{m}^{+}$ $\leq\max_{i\in[K]}\left(\|\mu_{i}^{+}-\mu_{i}^{*}\|\vee|\pi_{i}^{+}-\pi_{i}^{*}|/\pi_{i}^{*}\right)+\max_{i\in[K]}\left(\|\hat{\mu}_{i}^{+}-\mu_{i}^{+}\|\vee|\hat{\pi}_{i}^{+}-\pi_{i}^{+}|/\pi_{i}^{*}\right)$
$=D_{m}^{+}+\max_{i\in[K]}\left(\|\hat{\mu}_{i}^{+}-\mu_{i}^{+}\|\vee|\hat{\pi}_{i}^{+}-\pi_{i}^{+}|/\pi_{i}^{*}\right)$
We first claim that with probability at least $21-\delta/$ 2,
$\max_{i\in[K]}\left(\|\hat{\mu}_{i}^{+}-\mu_{i}^{+}\|\vee|\hat{\pi}_{i}^{+}-\pi_{i}^{+}|/\pi_{i}^{*}\right)\leq\tilde{c}_{2}\left(R_{\max}+d\left(1+\sqrt{\frac{2\log(\frac{2n}{\delta})}{d}}\right)\right)n\varepsilon.$ (19)
Then by Theorem C.5, with probability at least $1-\delta$, we have
$\hat{D}_{m}^{+}$ $\leq D_{m}^{+}+\max_{i\in[K]}\left(\|\hat{\mu}_{i}^{+}-\mu_{i}^{+}\|\vee|\hat{\pi}_{i}^{+}-\pi_{i}^{+}|/\pi_{i}^{*}\right)$
$\leq\frac{1}{2}+\frac{\tilde{c}}{(1-a)\pi_{min}}\sqrt{\frac{R_{max}(R_{max}\vee d)\log{\left(\frac{12K}{\delta}\right)}}{n}}+\tilde{c}_{2}\left(R_{\max}+d\left(1+\sqrt{\frac{2\log(\frac{2n}{\delta})}{d}}\right)\right)n\varepsilon$
$\leq a.$
Now we only need to prove Eq. 19. By the construction in Section C.2, we can see that $w_{\ell i}$ in first step can be well calculated, thus $|\hat{\pi}_{i}^{+}-\pi_{i}^{+}|=0$ and the error comes only from the calculation of ${\{\hat{\mu}_{i}^{+}\}}$. Recall that $\mu_{i}^{+}=\frac{\sum_{\ell=1}^{n}w_{\ell i}X_{\ell}}{\sum_{\ell=1}^{n}w_{\ell i}}$ and
$\hat{\mu}_{i}^{+}=\frac{\sum_{\ell=1}^{n}\exp{\left(\widehat{\log}({w}_{\ell i})\right)}X_{\ell}}{\sum_{l=1}^{n}\exp{\left(\widehat{\log}({w}_{\ell i})\right)}}.$
Recall that
$w_{\ell i}=\frac{\pi_{i}\exp(-\|X_{\ell}-\mu_{i}\|^{2}/2)}{\sum_{j=1}^{K}\pi_{j}\exp(-\|X_{\ell}-\mu_{j}\|^{2}/2)}=\frac{1}{1+\sum_{j\neq i}\frac{\pi_{j}}{\pi_{i}}\exp{\left({\left(\mu_{j}-\mu_{i}\right)}^{\top}X_{l}-\|\mu_{j}\|^{2}/2+\|\mu_{i}\|^{2}/2\right)}}.$
By the initial condition Eq. 10 and Eq. 9, we have
$\|\mu_{j}-\mu_{i}\|\leq R_{\max}+2*\frac{1}{16}R_{\min}=O(R_{\max}),~{}\|\mu_{j}\|^{2}=O(R_{\max}^{2}).$
[…]
$\sup_{\ell\in[n]}\|X_{\ell}\|\leq R_{\max}+d{\left(1+\sqrt{\frac{2\log(\frac{2n}{\delta})}{d}}\right)}=\tilde{O}(R_{\max}+d).$
Combine all things together, we get that with probability at least $21-\delta/$ 2,
$w_{\ell i}^{-1}\leq\exp{\left(\tilde{O}(K_{0}R_{\max}(R_{\max}+d_{0}))\right)},~{}\forall\ell\in[n]\text{ and }i\in[K].$
Thus taking $A=\exp{\left(\tilde{O}(K_{0}R_{\max}(R_{\max}+d_{0}))\right)}$ and and $\delta=\varepsilon$ in Lemma C.4, we can get $|\log-\widehat{\log}|\big{|}_{[1/A,A]}\leq\varepsilon$. Then by Lemma C.7, we have
$\|\hat{\mu}_{i}^{+}-\mu_{i}^{+}\|$ $=\left\|\frac{\sum_{\ell=1}^{n}\exp{\left(\widehat{\log}({w}_{\ell i})\right)}X_{\ell}}{\sum_{l=1}^{n}\exp{\left(\widehat{\log}({w}_{\ell i})\right)}}-\frac{\sum_{\ell=1}^{n}\exp{\left(\log w_{\ell i}\right)}X_{\ell}}{\sum_{\ell=1}^{n}\exp{\left(\log w_{\ell i}\right)}}\right\|$
$\leq\sum_{\ell=1}^{n}\left\|\frac{\exp{\left(\widehat{\log}({w}_{\ell i})\right)}X_{\ell}}{\sum_{l=1}^{n}\exp{\left(\widehat{\log}({w}_{\ell i})\right)}}-\frac{\exp{\left(\log w_{\ell i}\right)}X_{\ell}}{\sum_{\ell=1}^{n}\exp{\left(\log w_{\ell i}\right)}}\right\|$
$\leq\sup_{\ell\in[n]}\|X_{\ell}\|{\left(\sum_{\ell=1}^{n}\left|\frac{\exp{\left(\widehat{\log}({w}_{\ell i})\right)}}{\sum_{l=1}^{n}\exp{\left(\widehat{\log}({w}_{\ell i})\right)}}-\frac{\exp{\left(\log w_{\ell i}\right)}}{\sum_{\ell=1}^{n}\exp{\left(\log w_{\ell i}\right)}}\right|\right)}$
$\leq n{\left(R_{\max}+d+\sqrt{{2d\log\left(\frac{2n}{\delta}\right)}}\right)}{\left(\exp{\left(2\left\|{\left(\widehat{\log}({w}_{\ell i})\right)}_{\ell}-{\left({\log}({w}_{\ell i})\right)}_{\ell}\right\|_{\infty}\right)}-1\right)}$
$\leq 4n{\left(R_{\max}+d+\sqrt{{2d\log\left(\frac{2n}{\delta}\right)}}\right)}\varepsilon,~{}\forall i\in[K].$
Thus Eq. 19 is proved. The parameter bounds can be directly computed by the construction in Section C.2 and Lemma C.4.
Theorem C.7.
Fix $10<\delta,\beta<$ 1 and $11/2<a<$ 1. Let $C\geq 128$ be a large enough universal constant. Suppose the separation condition Eq. 8 holds and suppose the initialization parameter input to transformer satisfies ${D}_{m}\leq a$. If $n$ is suffcient large and $K_{0}\varepsilon\leq 1/100$ sufficient small such that
$\epsilon(n,\varepsilon,\delta,a):=\frac{\tilde{c}_{1}}{\left(1-a\right)\pi_{\min}}\sqrt{\frac{Kd\log(\frac{\tilde{C}n}{\delta})}{n}}+\tilde{c}_{2}{\left(\frac{1}{\pi_{\min}}+n{\left(R_{\max}+d+\sqrt{2d\log\left(\frac{2n}{\delta}\right)}\right)}\right)}\varepsilon<a(1-\beta),$
where $\tilde{c}_{1}$, $\tilde{c}_{2}$ are universal constants, $\tilde{C}=144K^{2}(\sqrt{d}+2R_{max}+\frac{1}{1-a})^{2}$. Then there exists a 2-layer transformer ${\rm TF}_{\boldsymbol{\Theta}}$ such that
$\hat{D}_{m}^{+}\leq\beta{D}_{m}+\epsilon(n,\varepsilon,\delta,a)\leq a$
uniformly holds with probability at least $1-\delta$. Moreover, ${\rm TF}_{\boldsymbol{\Theta}}$ falls within the class ${\mathcal{F}}$ with parameters satisfying:
$D=O(d_{0}+K_{0}),D^{\prime}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\varepsilon^{-1}\right)},M=O(1),\log B_{{\boldsymbol{\Theta}}}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\right)}.$
Proof.
Similar to the proof of Theorem C.6, using Theorem C.4, we only need to prove that with probability at least $21-\delta/$ 2,
$\max_{i\in[K]}\left(\|\hat{\mu}_{i}^{+}-\mu_{i}^{+}\|\vee|\hat{\pi}_{i}^{+}-\pi_{i}^{+}|/\pi_{i}^{*}\right)\leq\tilde{c}_{2}{\left(\frac{1}{\pi_{\min}}+n{\left(R_{\max}+d+\sqrt{2d\log\left(\frac{2n}{\delta}\right)}\right)}\right)}\varepsilon.$ (20)
Define $u_{\ell}=(u_{\ell,1},\cdots,u_{\ell,K})^{\top}$, $\hat{u}_{\ell}=(\hat{u}_{\ell,1},\cdots,\hat{u}_{\ell,K})^{\top}$, where $u_{\ell,i}=\log\pi_{i}+\mu_{i}^{\top}X_{\ell}-1/2\|\mu_{i}\|^{2}$ and $\hat{u}_{\ell,i}=\widehat{\log}\pi_{i}+\mu_{i}^{\top}X_{\ell}-1/2\widehat{\|\mu_{i}\|^{2}}$ By the construction in Section C.2 and Corollary C.1, we have
$\|\hat{{\mathbf{w}}}_{\ell}-{\mathbf{w}}_{\ell}\|_{\infty}=\left\|\frac{e^{\hat{u}_{\ell}}}{\|e^{\hat{u}_{\ell}}\|_{1}}-\frac{e^{u_{\ell}}}{\|e^{u_{\ell}}\|_{1}}\right\|_{\infty}\leq\exp{\left(2\|\hat{u}_{\ell}-u_{\ell}\|_{\infty}\right)}-1,~{}\forall\ell\in[n].$
Now taking $\delta=\varepsilon$, $A={\left((1-a)\pi_{\min}\right)}^{-1}$ in Lemma C.4 and $\delta=\varepsilon/K$, $A=(R_{\max}+a)^{2}$ in Lemma C.5, we have $\|\hat{u}_{\ell}-u_{\ell}\|_{\infty}\leq 3\varepsilon/2$, hence
$\|\hat{{\mathbf{w}}}_{\ell}-{\mathbf{w}}_{\ell}\|_{\infty}\leq\exp{\left(2\|\hat{u}_{\ell}-u_{\ell}\|_{\infty}\right)}-1\leq\exp(3\varepsilon)-1\leq 6\varepsilon,~{}\forall\ell\in[n].$
Then by the construction in Section C.2, we have
$|\hat{\pi}_{i}^{+}-\pi_{i}^{+}|\leq 6\varepsilon,~{}\forall i\in[K].$ (21)
For the term $\|\hat{\mu}_{i}^{+}-\mu_{i}^{+}\|$, we can calculate it similar to the proof of Theorem C.6. First, we recall that with probability at least $21-\delta/$ 2,
$w_{\ell i}^{-1}\leq\exp{\left(\tilde{O}(K_{0}R_{\max}(R_{\max}+d_{0}))\right)},~{}\forall\ell\in[n]\text{ and }i\in[K].$
Similarly, for $\hat{w}_{\ell,i}$, we can also get(just calculate again) that with probability at least $21-\delta/$ 2,
$\hat{w}_{\ell i}^{-1}\leq\exp{\left(\tilde{O}(K_{0}R_{\max}(R_{\max}+d_{0}))\right)},~{}\forall\ell\in[n]\text{ and }i\in[K].$
Then following the same argument in Theorem C.6, taking $A=\exp{\left(\tilde{O}(K_{0}R_{\max}(R_{\max}+d_{0}))\right)}$ and and $\delta=\varepsilon$ in Lemma C.4, we have also
$\|\hat{\mu}_{i}^{+}-\mu_{i}^{+}\|\leq 4n{\left(R_{\max}+d+\sqrt{{2d\log\left(\frac{2n}{\delta}\right)}}\right)}\varepsilon,~{}\forall i\in[K].$ (22)
Combining Eq. 21 and Eq. 22, Eq. 20 is proved. The parameter bounds can be directly computed by the construction in Section C.2, Lemma C.4, Lemma C.5 and the parameter $A$, $\delta$ taken in the proof.
C.4 Proof of Theorem C.1
First, by Theorem C.6 and the first condition in Theorem C.1 , there exist a $2$-layer transformer ${\rm TF}_{{\boldsymbol{\Theta}}_{1}}$ such that
$D_{{\boldsymbol{\Theta}}_{1}}^{\tt TF}\leq a,$ (23)
holds with probability at least $21-\delta/$ 2. Then using Theorem C.3, Eq. 23 and the second condition in Theorem C.1, there $2$-layer transformer ${\rm TF}_{{\boldsymbol{\Theta}}_{2}}$ such that
$D_{{\boldsymbol{\Theta}}_{1}\cup{\boldsymbol{\Theta}}_{2}}^{\tt TF}\leq\beta D_{{\boldsymbol{\Theta}}_{1}}^{\tt TF}+\epsilon(n,\varepsilon,\delta/2,a)\leq a,$
uniformly holds with probability at least $21-\delta/$ 2. Denote as ${\boldsymbol{\Theta}}_{2}^{L}=\cup_{\ell\in[L]}{\boldsymbol{\Theta}}_{2}$. Thus, for any $L\in\mathbb{N}$, by reduction, we have
$D_{{\boldsymbol{\Theta}}_{1}\cup{\boldsymbol{\Theta}}_{2}^{L}}^{\tt TF}\leq\beta^{L}D_{{\boldsymbol{\Theta}}_{1}}^{\tt TF}+{\left(1+\beta+\cdots+\beta^{L-1}\right)}\epsilon(n,\varepsilon,\delta/2,a),$
uniformly holds with probability at least $21-\delta/$ 2. Combine all things together, we have, for any $L\in\mathbb{N}$,
$D_{{\boldsymbol{\Theta}}_{1}\cup{\boldsymbol{\Theta}}_{2}^{L}}^{\tt TF}$ $\leq\beta^{L}D_{{\boldsymbol{\Theta}}_{1}}^{\tt TF}+{\left(1+\beta+\cdots+\beta^{L-1}\right)}\epsilon(n,\varepsilon,\delta/2,a)$
$\leq\beta^{L}a+\frac{1}{1-\beta}\epsilon(n,\varepsilon,\delta/2,a)$
holds with probability at least $1-\delta$ (Note that the definitions of $\epsilon{\left(\cdot\right)}$ in Theorem C.7 and Theorem C.1 differ slightly). The parameter bounds can be directly computed by Theorem C.6 and Theorem C.7. The theorem is proved.
Appendix D Formal statement of Theorem 2 and proofs
Following Appendix C, we implement $\operatorname{Readin}$ as an identity transformation and define $\operatorname{Readout}$ to extract targeted matrix elements hence they are both fixed functions.
D.1 Formal statement of Theorem 2
In this section, we give the formal statement of Theorem 2. First, we need to introduce the embeddings of the transformer. Let ${\mathbf{T}}$ be the matrix representation of the cubic tensor $T$, which is
${\mathbf{T}}\mathrel{\mathop{:}}=\begin{bmatrix}{\mathbf{t}}_{1},{\mathbf{t}}_{2},\cdots,{\mathbf{t}}_{d}\end{bmatrix}\mathrel{\mathop{:}}=\begin{bmatrix}T_{:,1,1}&T_{:,2,1}&\cdots&T_{:,d,1}\\ T_{:,1,2}&T_{:,2,2}&\cdots&T_{:,d,2}\\ \vdots&\vdots&\ddots&\vdots\\ T_{:,1,d}&T_{:,2,d}&\cdots&T_{:,d,d}\end{bmatrix}\in\mathbb{R}^{d^{2}\times d},$
where $T_{:,i,j}={\left(T_{1,i,j},T_{2,i,j},\cdots,T_{d,i,j}\right)}\in\mathbb{R}^{d}$, $i,j\in[d]$. For dimension adaptation, we assume $d\leq d_{0}$. The augment version of ${\mathbf{T}}$ is defined as
$\overline{{\mathbf{T}}}\mathrel{\mathop{:}}=\begin{bmatrix}\overline{{\mathbf{t}}}_{1},\overline{{\mathbf{t}}}_{2},\cdots,\overline{{\mathbf{t}}}_{d_{0}}\end{bmatrix}\mathrel{\mathop{:}}=\begin{bmatrix}\overline{T}_{:,1,1}&\overline{T}_{:,2,1}&\cdots&\overline{T}_{:,d_{0},1}\\ \overline{T}_{:,1,2}&\overline{T}_{:,2,2}&\cdots&\overline{T}_{:,d_{0},2}\\ \vdots&\vdots&\ddots&\vdots\\ \overline{T}_{:,1,d_{0}}&\overline{T}_{:,2,d_{0}}&\cdots&\overline{T}_{:,d_{0},d_{0}}\end{bmatrix}\in\mathbb{R}^{d_{0}^{2}\times d_{0}},$ (24)
where $\overline{T}_{:,i,j}\in\mathbb{R}^{d_{0}}$. If $i\leq d$ and $j\leq d$, $\overline{T}_{:,i,j}={\left[T_{:,i,j}^{\top},{\mathbf{0}}_{d_{0}-d}^{\top}\right]}^{\top}$; Else $\overline{T}_{:,i,j}={\mathbf{0}}_{d_{0}}$. We construct the following input sequence:
${\mathbf{H}}=\begin{bmatrix}\overline{{\mathbf{t}}}_{1}&\overline{{\mathbf{t}}}_{2}&\dots&\overline{{\mathbf{t}}}_{d}\\ {\mathbf{p}}_{1}&{\mathbf{p}}_{2}&\dots&{\mathbf{p}}_{d}\end{bmatrix}\in\mathbb{R}^{D\times d},~{}{\mathbf{p}}_{i}=\begin{bmatrix}\overline{v}^{(0)}\\ {\mathbf{e}}_{i}\\ 1\\ d\\ {\mathbf{0}}_{\tilde{D}}\end{bmatrix},$ (25)
where $\overline{{\mathbf{t}}}_{i}\in\mathbb{R}^{d_{0}^{2}}$ is defined as Eq. 24, $\overline{v}^{(0)}={\left[v^{(0)\top},{\mathbf{0}}_{d_{0}-d}^{\top}\right]}^{\top}\in\mathbb{R}^{d_{0}}$, ${\mathbf{e}}_{i}\in\mathbb{R}^{d_{0}}$ denotes the $i$-th standard unit vector in $\mathbb{R}^{d_{0}}$. We choose $D=O(d_{0}^{2})$ and $\tilde{D}=D-d_{0}^{2}-2d_{0}-2$ to get the encoding above. Then we give a rigorous definition of ReLU-activated transformer following [4].
Definition 4 (ReLU-attention layer).
A (self-)attention layer activated by ReLU function with $M$ heads is denoted as ${\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}}(\cdot)$ with parameters ${\boldsymbol{\Theta}}_{\tt attn}={\{({\mathbf{V}}_{m},{\mathbf{Q}}_{m},{\mathbf{K}}_{m})\}}_{m\in[M]}\subset\mathbb{R}^{D\times D}$. On any input sequence ${\mathbf{H}}\in\mathbb{R}^{D\times N}$,
$\widetilde{{\mathbf{H}}}={\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}}^{R}({\mathbf{H}})\mathrel{\mathop{:}}={\mathbf{H}}+\frac{1}{N}\sum_{m=1}^{M}({\mathbf{V}}_{m}{\mathbf{H}})\sigma{\left(({\mathbf{K}}_{m}{\mathbf{H}})^{\top}({\mathbf{Q}}_{m}{\mathbf{H}})\right)}\in\mathbb{R}^{D\times N},$
In vector form,
$\widetilde{{\mathbf{h}}}_{i}={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}}^{R}({\mathbf{H}})\right]}_{i}={\mathbf{h}}_{i}+\sum_{m=1}^{M}\frac{1}{N}\sum_{j=1}^{N}\sigma{\left({\left({\mathbf{Q}}_{m}{\mathbf{h}}_{i}\right)}^{\top}{\left({\mathbf{K}}_{m}{\mathbf{h}}_{j}\right)}\right)}{\mathbf{V}}_{m}{\mathbf{h}}_{j}.$
Here $\sigma(x)=x\vee 0$ denotes the ReLU function.
The MLP layer is the same as Definition 2. The ReLU-activated transformer is defined as follows.
Definition 5 (ReLU-activated transformer).
An $L$-layer transformer, denoted as ${\rm TF}_{{\boldsymbol{\Theta}}}^{R}(\cdot)$, is a composition of $L$ ReLU-attention layers each followed by an MLP layer:
${\rm TF}_{{\boldsymbol{\Theta}}}^{R}({\mathbf{H}})=\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(L)}}{\left({\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(L)}}^{R}{\left(\cdots\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(1)}}{\left({\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(1)}}^{R}{\left({\mathbf{H}}\right)}\right)}\right)}\right)}.$
Above, the parameter ${\boldsymbol{\Theta}}=({\boldsymbol{\Theta}}_{\tt attn}^{(1:L)},{\boldsymbol{\Theta}}_{\tt mlp}^{(1:L)})$ consists of the attention layers ${\boldsymbol{\Theta}}_{\tt attn}^{(\ell)}={\{({\mathbf{V}}^{(\ell)}_{m},{\mathbf{Q}}^{(\ell)}_{m},{\mathbf{K}}^{(\ell)}_{m})\}}_{m\in[M^{(\ell)}]}\subset\mathbb{R}^{D\times D}$ and the MLP layers ${\boldsymbol{\Theta}}_{\tt mlp}^{(\ell)}=({\mathbf{W}}^{(\ell)}_{1},{\mathbf{W}}^{(\ell)}_{2})\in\mathbb{R}^{D^{(\ell)}\times D}\times\mathbb{R}^{D\times D^{(\ell)}}$.
Similar to Section 4.1, We consider the following function class of transformer.
${\mathcal{F}}\mathrel{\mathop{:}}={\mathcal{F}}(L,D,D^{\prime},M,B_{\boldsymbol{\Theta}})={\left\{{\rm TF}_{{\boldsymbol{\Theta}}}^{R},{\left|\!\left|\!\left|{{\boldsymbol{\Theta}}}\right|\!\right|\!\right|}\leq B_{\boldsymbol{\Theta}},D^{(\ell)}\leq D^{\prime},M^{\ell}\leq M,\ell\in[L]\right\}}.$
Now we can give the formal statement of Theorem 2.
Theorem D.8 (Formal version of Theorem 2).
There exists a transformer ${\rm TF}_{\boldsymbol{\Theta}}$ with ReLU activation such that for any $d\leq d_{0}$, $T\in\mathbb{R}^{d\times d\times d}$ and $v^{(0)}\in\mathbb{R}^{d}$, given the encoding Eq. 25, ${\rm TF}_{\boldsymbol{\Theta}}$ implements $L$ steps of Eq. 4 exactly. Moreover, ${\rm TF}_{\boldsymbol{\Theta}}$ falls within the class ${\mathcal{F}}$ with parameters satisfying:
$D=D^{\prime}=O(d_{0}^{2}),M=O(d_{0}),\log B_{{\boldsymbol{\Theta}}}\leq O(1).$
Remark D.1.
In fact, Theorem D.8 is also hold for attention-only transformers since the MLP layer do not use in the proof. To do that, we only need to add another head in every odd attention layer to clean the terms ${\{d\overline{v}_{i}\}}$. For details, see the proof.
Remark D.2.
Readers might question why the normalization step is omitted in our theorem. The key challenge is that we have absolutely no knowledge of a lower bound for $\left\|{T{\left(I,v^{(j)},v^{(j)}\right)}}\right\|$. Without this bound, approximating the normalization step becomes infeasible.
Remark D.3.
The use of the ReLU activation function here is primarily for technical convenience and does not alter the fundamental nature of the attention mechanism. Several studies have demonstrated that transformers with ReLU-based attention perform comparably to those using softmax attention[41, 4, 16].
D.2 Proof of Theorem D.8
Proof.
For simplicity, we only proof the case that $\sigma(x)=x$ in the attention layer. For ReLU activated transformer, the result can be similarly proved by $\operatorname{ReLU}(x)-\operatorname{ReLU}(-x)=x$ and the $\sigma(x)=x$ case. Hence we omit the notation $\sigma$ in the following proof. We take ${\mathbf{H}}^{(0)}={\mathbf{H}}$. In the first attention layer, consider the following attention structures:
${\mathbf{Q}}^{(1)}{\mathbf{h}}_{i}^{(0)}=\begin{bmatrix}{\mathbf{e}}_{i}\\ {\mathbf{0}}\end{bmatrix},~{}{\mathbf{K}}^{(1)}{\mathbf{h}}_{j}^{(0)}=\begin{bmatrix}\overline{v}^{(0)}\\ {\mathbf{0}}\end{bmatrix},~{}{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{(0)}=\begin{bmatrix}{\mathbf{0}}_{d_{0}^{2}}\\ {\mathbf{0}}_{d_{0}}\\ 0\\ 0\\ d\\ {\mathbf{0}}\end{bmatrix}.$
After the attention operation, we have
$\widetilde{{\mathbf{h}}}_{i}^{(1)}={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(1)}}^{R}{\left({\mathbf{H}}^{0}\right)}\right]}_{:,i}={\mathbf{h}}_{i}^{(0)}+\frac{1}{d}\sum_{j=1}^{d}{\left({\left({\mathbf{Q}}^{(1)}{\mathbf{h}}_{i}^{0}\right)}^{\top}{\left({\mathbf{K}}^{(1)}{\mathbf{h}}_{j}^{0}\right)}\right)}{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{0}={\mathbf{h}}_{i}^{(0)}+\begin{bmatrix}{\mathbf{0}}_{d_{0}^{2}}\\ {\mathbf{0}}_{d_{0}}\\ 0\\ 0\\ d\overline{v}^{(0)}_{i}\\ {\mathbf{0}}\end{bmatrix}=\begin{bmatrix}\overline{{\mathbf{t}}}_{i}\\ \overline{v}^{(0)}\\ 1\\ d\\ d\overline{v}^{(0)}_{i}\\ {\mathbf{0}}\end{bmatrix},~{}i\in[d].$
Then we use a two-layer MLP to implement identity operation, which is
${\mathbf{h}}_{i}^{(1)}=\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(1)}}{\left(\widetilde{{\mathbf{h}}}_{i}^{(1)}\right)}=\begin{bmatrix}\overline{{\mathbf{t}}}_{i}\\ \overline{v}^{(0)}\\ 1\\ d\\ d\overline{v}^{(0)}_{i}\\ {\mathbf{0}}\end{bmatrix},~{}i\in[d].$
Now we use an attention layer with $d_{0}+1$ heads to implement the power iteration step of the cubic tensor. Consider the following attention structure:
${\mathbf{Q}}_{m}^{(2)}{\mathbf{h}}_{i}^{(1)}=\begin{bmatrix}\overline{v}^{(0)}_{m}\\ {\mathbf{0}}\end{bmatrix},~{}{\mathbf{K}}_{m}^{(2)}{\mathbf{h}}_{j}^{(1)}=\begin{bmatrix}d\overline{v}^{(0)}_{j}\\ {\mathbf{0}}\end{bmatrix},~{}{\mathbf{V}}_{m}^{(2)}{\mathbf{h}}_{j}^{(1)}=\begin{bmatrix}{\mathbf{0}}_{d_{0}^{2}}\\ \overline{T}_{:,j,m}\\ {\mathbf{0}}\end{bmatrix},~{}m\in[d_{0}],$
and
${\mathbf{Q}}_{d_{0}+1}^{(2)}{\mathbf{h}}_{i}^{(1)}=\begin{bmatrix}1\\ {\mathbf{0}}\end{bmatrix},~{}{\mathbf{K}}_{d_{0}+1}^{(2)}{\mathbf{h}}_{j}^{(1)}=\begin{bmatrix}d\\ {\mathbf{0}}\end{bmatrix},~{}{\mathbf{V}}_{d_{0}+1}^{(2)}{\mathbf{h}}_{j}^{(1)}=\begin{bmatrix}{\mathbf{0}}_{d_{0}^{2}}\\ -\overline{v}^{(0)}\\ {\mathbf{0}}\end{bmatrix}.$
After the attention operation, we have
$\widetilde{{\mathbf{h}}}_{i}^{(2)}$ $={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(2)}}^{R}{\left({\mathbf{H}}^{(1)}\right)}\right]}_{:,i}$
$={\mathbf{h}}_{i}^{(1)}+\sum_{m=1}^{d_{0}}\frac{1}{d}\sum_{j=1}^{d}{\left({\left({\mathbf{Q}}^{(2)}_{m}{\mathbf{h}}_{i}^{(1)}\right)}^{\top}{\left({\mathbf{K}}^{(2)}_{m}{\mathbf{h}}_{j}^{(1)}\right)}\right)}{\mathbf{V}}^{(2)}_{m}{\mathbf{h}}_{j}^{(1)}+\frac{1}{d}\sum_{j=1}^{d}{\left({\left({\mathbf{Q}}^{(2)}_{d_{0}+1}{\mathbf{h}}_{i}^{(1)}\right)}^{\top}{\left({\mathbf{K}}^{(2)}_{d_{0}+1}{\mathbf{h}}_{j}^{(1)}\right)}\right)}{\mathbf{V}}^{(2)}_{d_{0}+1}{\mathbf{h}}_{j}^{(1)}$
$={\mathbf{h}}_{i}^{(1)}+\sum_{m=1}^{d_{0}}\frac{1}{d}\sum_{j=1}^{d}{\left(d\overline{v}^{(0)}_{m}\overline{v}^{(0)}_{j}\right)}\begin{bmatrix}{\mathbf{0}}_{d_{0}^{2}}\\ \overline{T}_{:,j,m}\\ {\mathbf{0}}\end{bmatrix}+\frac{1}{d}\sum_{j=1}^{d}d\begin{bmatrix}{\mathbf{0}}_{d_{0}^{2}}\\ -\overline{v}^{(0)}\\ {\mathbf{0}}\end{bmatrix}$
$={\mathbf{h}}_{i}^{(1)}+\begin{bmatrix}{\mathbf{0}}_{d_{0}^{2}}\\ \overline{v}^{(1)}\\ {\mathbf{0}}\end{bmatrix}+\begin{bmatrix}{\mathbf{0}}_{d_{0}^{2}}\\ -\overline{v}^{(0)}\\ {\mathbf{0}}\end{bmatrix}$
$=\begin{bmatrix}\overline{{\mathbf{t}}}_{i}\\ \overline{v}^{(1)}\\ 1\\ d\\ d\overline{v}^{(0)}_{i}\\ {\mathbf{0}}\end{bmatrix},~{}i\in[d],$
where $\overline{v}^{(1)}={\left[{v}^{(1)\top},{\mathbf{0}}_{d_{0}-d}^{\top}\right]}^{\top}$ and ${v}^{(1)}=\sum_{j,m\in[d]}v_{m}^{(0)}v_{j}^{(0)}T_{:,j,m}$.
Then we use a two-layer MLP to clean the term $d\overline{v}^{(0)}_{i}$, which is
${\mathbf{h}}_{i}^{(2)}=\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(2)}}{\left(\widetilde{{\mathbf{h}}}_{i}^{(2)}\right)}=\begin{bmatrix}\overline{{\mathbf{t}}}_{i}\\ \overline{v}^{(1)}\\ 1\\ d\\ {\mathbf{0}}\end{bmatrix},~{}i\in[d].$
Similarly, for any $\ell\in\mathbb{N}_{+}$, we have
${\mathbf{h}}_{i}^{(2\ell)}={\left[\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(2)}}{\left({\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(2)}}{\left(\mathrm{MLP}_{{\boldsymbol{\Theta}}_{\tt mlp}^{(1)}}{\left({\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(1)}}{\left({\mathbf{H}}^{(2\ell-2)}\right)}\right)}\right)}\right)}\right]}_{:,i}=\begin{bmatrix}\overline{{\mathbf{t}}}_{i}\\ \overline{v}^{(\ell)}\\ 1\\ d\\ {\mathbf{0}}\end{bmatrix},~{}i\in[d].$
The parameter bounds can be directly computed by the construction above. The theorem is proved.
Appendix E More on empirical studies
E.1 More on experimental setups
Anisotropic adjustments
We consider anisotropic Gaussian mixtures that takes the following form: A $K$-component anisotropic Gaussian mixture distribution is defined with parameters ${\boldsymbol{\theta}}={\boldsymbol{\pi}}\cup{\boldsymbol{\mu}}\cup{\boldsymbol{\sigma}}$, where ${\boldsymbol{\pi}}\mathrel{\mathop{:}}={\{\pi_{1}^{\ast},\pi_{2}^{\ast},\cdots,\pi_{K}^{\ast}\}}$, $\pi_{k}^{\ast}\in\mathbb{R}$, ${\boldsymbol{\mu}}={\{\mu_{1}^{\ast},\mu_{2}^{\ast},\cdots,\mu_{K}^{\ast}\}},\mu_{k}^{\ast}\in\mathbb{R}^{d}$, $k\in[K]$ and ${\boldsymbol{\sigma}}={\{\sigma_{1}^{\ast},\sigma_{2}^{\ast},\cdots,\sigma_{K}^{\ast}\}},\sigma_{k}^{\ast}\in\mathbb{R}_{+}^{d}$, $k\in[K]$. A sample $X_{i}$ from the aforementioned anisotropic GMM is expressed as:
$X_{i}=\mu_{y_{i}}^{\ast}+\sigma_{y_{i}}^{\ast}Z_{i},$ (26)
where ${\{y_{i}\}}_{i\in[N]}$ are iid discrete random variables with ${\mathbb{P}}{\left(y=k\right)}=\pi_{k}^{\ast}$ for $k\in[K]$ and ${\{Z_{i}\}}_{i\in[N]}$ are iid standard Gaussian random vector in $\mathbb{R}^{d}$. Analogous to that in the isotropic case and overload some notations, we define an anisotropic GMM task to be $\mathcal{T}=({\mathbf{X}},{\boldsymbol{\theta}},K)$.
To adapt the TGMM framework to be compatible to anisotropic problems, we expand the output dimension of the attentive pooling module from $(d+K)\times K$ to $(d+2K)\times K$, with the additional $K$ rows reserved for the estimate $\widehat{{\boldsymbol{\sigma}}}$ of ${\boldsymbol{\sigma}}$, with the corresponding estimation loss function augmented with a scale estimation part:
$\widehat{L}_{n}{\left({\boldsymbol{\Theta}}\right)}=\frac{1}{n}\sum_{i=1}^{n}\ell_{\mu}(\widehat{{\boldsymbol{\mu}}}_{i},{\boldsymbol{\mu}}_{i})+\ell_{\pi}(\widehat{{\boldsymbol{\pi}}}_{i},{\boldsymbol{\pi}}_{i})+\ell_{\sigma}(\widehat{{\boldsymbol{\sigma}}}_{i},{\boldsymbol{\sigma}}_{i}),$ (27)
where the loss function $\ell_{\sigma}$ is chosen as the mean-square loss. During the experiments, we inherit configurations from those of isotropic counterparts, except for the calculation of the $\ell_{2}$-error metric, where we additionally considered contributions from the estimation error of scales.
Configurations related to Mamba2 architecture
We adopt a Mamba2 [8] model comprising $12$-layers and $128$-dimensional hidden states, with the rest hyper-parameters chosen so as to approximately match the number of a $12$-layer transformer with $128$-dimensional hidden states. As the Mamba series of models are essentially recurrent neural networks (RNNs), we tested two different kinds of $\operatorname{Readout}$ design with either (i). the attentive pooling module as used in the case of transformer backbone and (ii). a more natural choice of using simply the last hidden state to decode all the estimates, as RNNs compress input information in an ordered fashion. We observe from our empirical investigations that using attentive pooling yields better performance even with a Mamba2 backbone. The other training configurations are cloned from those in TGMM experiments with transformer backbones.
Software and hardware infrastructures
Our framework is built upon PyTorch [37] and transformers [46] libraries, which are open-source software released under BSD-style  [^2] https://github.com/pytorch/pytorch/blob/master/LICENSE and Apache license  [^3] https://github.com/huggingface/transformers/blob/main/LICENSE. The code implementations will be open-sourced after the reviewing process of this paper. All the experiments are conducted using $8$ NVIDIA A100 GPUs with $80$ GB memory each.
E.2 A complete report regarding different evaluation metrics
In this section, we present complete reports of empirical performance regarding the evaluation problems mentioned in section 3. Aside from the $\ell_{2}$-error metric that was reported in section 3.2, we additionally calculated all the experimental performance under the following metrics:
Clustering accuracy
We compare estimated cluster membership with the true component assignment, after adjusting for permutation invariance as mentioned in section 2.3.
Log-likelihood
We compute average log-likelihood as a standard metric in unsupervised statistical estimation.
The results are reported in figure 10, 11, 12, 13 and 14, respectively. According to the evaluations, the learned TGMM models show comparable clustering accuracy against the spectral algorithm and outperform EM algorithm when $K>2$ across all comparisons. Regarding the log-likelihood metric, TGMM demonstrates comparable performance with the other two classical algorithms in comparatively lower dimensional cases. i.e., $d\in{\{2,8\}}$, but underperforms both baselines in larger dimensional problems. We conjecture that is might be due to the fact that EM algorithm is essentially a maximum-likelihood algorithm [10], while the TGMM estimation objective (2) is not explicitly related to likelihood-based training.
(a) $\ell_{2}$-error
(b) clustering accuracy
(c) log-likelihood
Figure 10: Performance comparison between TGMM and two classical algorithms, reported in three metrics.
(a) $\ell_{2}$-error
(b) clustering accuracy
(c) log-likelihood
Figure 11: Assessments of TGMM under test-time task distribution shifts I: A line with $N_{0}^{\text{train}}\rightarrow N^{\text{test}}$ draws the performance of a TGMM model trained over tasks with sample size randomly sampled in $[N_{0}^{\text{train}}/2,N_{0}^{\text{train}}]$ and evaluated over tasks with sample size $N^{\text{test}}$. We can view the configuration $128\rightarrow 128$ as an in-distribution test and rest as out-of-distribution tests.
(a) $\ell_{2}$-error
(b) clustering accuracy
(c) log-likelihood
Figure 12: Assessments of TGMM under test-time task distribution shifts II: $\ell_{2}$-error of estimation when the test-time tasks $\mathcal{T}^{\text{test}}$ are sampled using a mean vector sampling distribution $p_{\mu}^{\text{test}}$ different from the one used during training.
(a) $\ell_{2}$-error
(b) clustering accuracy
(c) log-likelihood
Figure 13: Performance comparisons between TGMM using transformer and Mamba2 as backbone, reported in three metrics.
(a) $\ell_{2}$-error
(b) clustering accuracy
(c) log-likelihood
Figure 14: Performance comparison between TGMM and the EM algorithm on anisotropic GMM tasks, reported in three metrics
E.3 On the impact of inference-time sample size $N$
Motivated by the classical statistical phenomenon that estimation quality tends to improve with sample size, we test whether TGMM’s estimation performance increases as $N$ goes up. We run corresponding experiments by varying the sample size to be $N\in{\{32,64,128\}}$ during both train and inference, while controlling other experimental configurations same as those in section 3.1. The results are reported in $\ell_{2}$-error, clustering accuracy as log-likelihood and summarized in figure 15. The results exhibit a clear trend that aligns with our hypothesis, justifying the TGMM learning process as learning a statistically meaningful algorithm for solving GMMs.
(a) $\ell_{2}$-error
(b) clustering accuracy
(c) log-likelihood
Figure 15: Performance comparison between TGMM models trained under varying configurations of sample-size. For example, $N=64$ means that the model is trained over GMM tasks with (randomly chosen) sample sizes within the range $[32,64]$ and tested on tasks with sample size $64$.
E.4 On the impact of backbone scale
The scaling phenomenon is among the mostly discussed topics in modern AI, as choosing a suitable scale is often critical to the performance of transformer-based architectures like LLMs. In this section we investigate the scaling properties of TGMM via comparing performances produced by varying sizes of backbones that differ either in per-layer width (i.e., the dimension of attention embeddings) or in the total number of layers $L$. With the rest hyper-parameters controlled to be the same as those in section 3.1. The results are reported in three metrics and summarized in figure 16 and figure 17, respectively. According to these investigations, while in general a larger-sized backbone yields slightly better performance as compared to smaller ones. The performance gaps remain mild especially for tasks with relative lower complexity, i.e., $K=2$. Consequently, even a $3$-layer transformer backbone is able to achieve non-trivial learning performance for solving isotropic GMMs, a phenomenon that was also observed in a recent work [17].
(a) $\ell_{2}$-error
(b) clustering accuracy
(c) log-likelihood
Figure 16: Performance comparison between TGMM under backbones of varying scales I: We fix embedding size at $d=128$ and tested over different number of transformer layers $L\in\{3,6,12\}$. Results are reported in three metrics.
(a) $\ell_{2}$-error
(b) clustering accuracy
(c) log-likelihood
Figure 17: Performance comparison between TGMM under backbones of varying scales II: We fix the number of transformer layers at $L=12$ and tested over different number of hidden states $d\in\{128,256,512\}$. Results are reported in three metrics.